### v2

# The Seifert-van Kampen Theorem via Computational Paths: A Formalized Approach to Computing Fundamental Groups

Abstract.
The Seifert-van Kampen theorem computes the fundamental group of a space from the fundamental groups of its constituents. We formalize this theorem within the framework of computational paths—an approach to equality where witnesses are explicit sequences of rewrites governed by the confluent, terminating $\mathrm{LND}_{\mathrm{EQ}}$-TRS.
Our contributions are: (i) pushouts as higher-inductive types with explicit path constructors; (ii) free products and amalgamated free products as quotients of word representations; (iii) an encode-decode proof establishing $\pi_{1}(\mathrm{Pushout}(A,B,C))\simeq\pi_{1}(A)\ast_{\pi_{1}(C)}\pi_{1}(B)$; and (iv) applications to the figure-eight ($\pi_{1}(S^{1}\vee S^{1})\simeq\mathbb{Z}\ast\mathbb{Z}$), 2-sphere ($\pi_{1}(S^{2})\simeq 1$), and 3-sphere ($\pi_{1}(S^{3})\simeq 1$ with Hopf fibration context).
Recent extensions include: higher homotopy groups $\pi_{n}$ via the weak $\omega$-groupoid structure (with $\pi_{2}$ abelian via Eckmann-Hilton, and $\pi_{2}\cong 1$ in the 1-groupoid truncated setting); truncation levels connecting the framework to HoTT n-types; automated path simplification tactics; basic covering space theory with $\pi_{1}$-actions on fibers; fibration theory with long exact sequences; and Eilenberg-MacLane space characterization ($S^{1}=K(\mathbb{Z},1)$).
The framework makes coherence witnesses explicit as rewrite derivations. The development is formalized in Lean 4 with over 25,000 lines of mechanized proofs. This demonstrates that the encode-decode method for higher-inductive types becomes fully constructive when path equality is decidable via normalization.
Key words and phrases:
Seifert-van Kampen theorem, computational paths, fundamental group, higher-inductive types, pushouts, free products, type theory
2020 Mathematics Subject Classification:
Primary: 55Q05; Secondary: 03B38, 18N10, 68V15
1. Introduction
The Seifert-van Kampen theorem, first proved by Herbert Seifert [16] and Egbert van Kampen [18], is one of the most powerful tools in algebraic topology for computing fundamental groups (see [7, 11] for modern treatments). It states that if a path-connected space $X$ is the union of two path-connected open subspaces $U$ and $V$ with path-connected intersection $U\cap V$, then the fundamental group $\pi_{1}(X)$ can be computed as the amalgamated free product:
$\pi_{1}(X)\cong\pi_{1}(U)\ast_{\pi_{1}(U\cap V)}\pi_{1}(V)$
In homotopy type theory (HoTT) [17], this theorem takes a particularly elegant form when expressed in terms of pushouts—a higher-inductive type (HIT) that generalizes the notion of gluing spaces together. The HoTT version of the Seifert-van Kampen theorem was developed by Favonia and Shulman [8], providing a synthetic proof that works entirely within type theory.
The present paper develops the Seifert-van Kampen theorem within the framework of computational paths [2, 13]—an alternative approach to equality in type theory where witnesses of equality are explicit sequences of rewrites. This approach offers several advantages:
  1. (1)
    Explicit witnesses: Every equality proof carries the full computational content showing why two terms are equal.
  2. (2)
    Decidable equality on paths: The rewrite equality relation ($\sim$) is decidable via normalization, making path equality computationally tractable.
  3. (3)
    Concrete coherence: Higher coherence witnesses (pentagon, triangle, etc.) are explicit rewrite derivations rather than abstract existence proofs.
Relation to Classical SVK.
The classical theorem requires $X=U\cup V$ with $U,V,U\cap V$ open and path-connected. The HIT version generalizes this: the pushout $\mathrm{Pushout}(A,B,C,f,g)$ corresponds to the union where $f$ and $g$ specify how the “intersection” $C$ embeds into both spaces. Path-connectedness assumptions are needed to ensure the fundamental group is well-defined; in the HIT setting, this becomes the assumption that all types are path-connected.
1.1. Main Contributions
This paper provides:
  1. (1)
    A formal definition of pushouts as higher-inductive types within the computational paths framework, including:
    • Point constructors $\mathrm{inl}:A\to\mathrm{Pushout}$ and $\mathrm{inr}:B\to\mathrm{Pushout}$
    • Path constructor $\mathrm{glue}:\prod_{c:C}\mathrm{Path}(\mathrm{inl}(f(c)),\mathrm{inr}(g(c)))$
    • Recursion and induction principles with computation rules
  2. (2)
    The construction of free products $G_{1}\ast G_{2}$ and amalgamated free products $G_{1}\ast_{H}G_{2}$ as quotients of word representations.
  3. (3)
    A complete encode-decode proof of the Seifert-van Kampen theorem:
    $\pi_{1}(\mathrm{Pushout}(A,B,C,f,g),\mathrm{inl}(f(c_{0})))\simeq\pi_{1}(A,f(c_{0}))\ast_{\pi_{1}(C,c_{0})}\pi_{1}(B,g(c_{0}))$
  4. (4)
    Applications:
    • The wedge sum: $\pi_{1}(A\vee B)\simeq\pi_{1}(A)\ast\pi_{1}(B)$
    • The figure-eight: $\pi_{1}(S^{1}\vee S^{1})\simeq\mathbb{Z}\ast\mathbb{Z}$
    • The 2-sphere: $\pi_{1}(S^{2})\simeq 1$ (trivial group)
    • The 3-sphere: $\pi_{1}(S^{3})\simeq 1$ with Hopf fibration context
  5. (5)
    A complete Lean 4 formalization available at [15].
1.2. Related Work
The Seifert-van Kampen theorem in HoTT was first formalized by Favonia and Shulman [8], who developed a general framework for computing fundamental groups of HITs. Their work uses the encode-decode method pioneered by Licata and Shulman [10] for computing $\pi_{1}(S^{1})\cong\mathbb{Z}$.
The computational paths approach to equality originates in work by de Queiroz and colleagues [2, 13, 14], building on earlier ideas about equality proofs as sequences of rewrites [3]. Our previous work established that computational paths form a weak groupoid [5] and can be used to calculate fundamental groups [4, 6].
The connection between rewriting and coherence has been developed by Kraus and von Raumer [9], who showed that confluence and termination suffice to build coherent higher-dimensional structure. Their key insight is that in a confluent, terminating rewrite system, any two parallel reductions to normal form can be connected by a canonical “diamond” of rewrite paths—and this extends to higher dimensions. For our $\mathrm{LND}_{\mathrm{EQ}}$-TRS, this means that higher coherences (such as the pentagon identity for associativity and the naturality conditions for transport) are automatically satisfied once we establish that paths with the same normal form are $\sim$-equivalent. This principle is crucial for the SVK proof, where glue naturality and amalgamation compatibility require reasoning about 2-dimensional path equalities.
1.3. Reading Guide
Readers interested in the mathematical content can focus on Sections 36, treating Lean code as pseudocode. Those interested in formalization should additionally consult Section 7 and the supplementary code. Key definitions are boxed for quick reference. We use the figure-eight space $S^{1}\vee S^{1}$ as a running example throughout.
1.4. Structure of the Paper
Section 2 reviews computational paths, rewrite equality, and fundamental groups. Section 3 defines pushouts as higher-inductive types. Section 4 constructs free products and amalgamated free products. Section 5 proves the Seifert-van Kampen theorem using the encode-decode method. Section 6 applies the theorem to compute fundamental groups of the figure-eight, 2-sphere, and 3-sphere (with Hopf fibration context). Section 7 discusses the Lean 4 formalization. Section 8 concludes.
2. Background: Computational Paths
2.1. Computational Paths
A computational path $p:\mathrm{Path}(a,b)$ from $a$ to $b$ (both terms of type $A$) is an explicit sequence of rewrite steps witnessing the equality $a=b$. The fundamental operations are:
  • Reflexivity: $\rho(a):\mathrm{Path}(a,a)$ — the empty sequence of rewrites.
  • Symmetry: If $p:\mathrm{Path}(a,b)$, then $\sigma(p):\mathrm{Path}(b,a)$ — reverse and invert each step.
  • Transitivity: If $p:\mathrm{Path}(a,b)$ and $q:\mathrm{Path}(b,c)$, then $p\cdot q:\mathrm{Path}(a,c)$ — concatenate the sequences.
  • Congruence: If $p:\mathrm{Path}(a,b)$ and $f:A\to B$, then $f_{*}(p):\mathrm{Path}(f(a),f(b))$.
  • Transport: If $p:\mathrm{Path}(a,b)$ and $P:A\to\mathrm{Type}$, then $\mathrm{transport}(P,p):P(a)\to P(b)$.
Notation.
Throughout this paper, we use:
  • $p\cdot q$ for path composition (transitivity)
  • $p^{-1}$ for path inverse (symmetry)
  • $f_{*}(p)$ for $\mathrm{congrArg}(f,p)$
  • $\rho(a)$ or simply $\rho$ for reflexivity
In Lean, a path is represented as a wrapper around an equality proof:
1structure Path {A : Type u} (a b : A) where
2 toEq : a = b -- the underlying equality proof
Remark 2.1 (Implementation Note).
The Lean implementation stores only the underlying equality proof, not an explicit rewrite sequence. The “rewrite sequence” interpretation is recovered via the Step and RwEq relations, which classify paths syntactically. This design choice simplifies the formalization while preserving the key property that path equality is captured by rewrite equivalence.
2.2. The Step Relation
The foundation of the computational paths framework is the Step relation, which defines single-step rewrites between paths. The $\mathrm{LND}_{\mathrm{EQ}}$-TRS consists of over 50 rewrite rules organized into categories:
  • Groupoid laws: $\rho^{-1}\rightsquigarrow\rho$, $(p^{-1})^{-1}\rightsquigarrow p$, $\rho\cdot p\rightsquigarrow p$, $\cdot\rho\rightsquigarrow p, pp$$\cdot p^{-1}\rightsquigarrow\rho$, $p^{-1}\cdot p\rightsquigarrow\rho$, $(p\cdot q)^{-1}\rightsquigarrow q^{-1}\cdot p^{-1}$, and $(p\cdot q)\cdot r\rightsquigarrow p\cdot(q\cdot r)$.
  • Type-specific rules: $\beta$-rules for products, sums, and functions; $\eta$-rules; transport laws.
  • Context rules: Allow rewrites inside larger expressions: if $p\rightsquigarrow q$ then $C[p]\rightsquigarrow C[q]$.
  • Canonicalization: $p\rightsquigarrow\mathrm{ofEq}(p.\mathrm{toEq})$ — any path can reduce to its canonical form.
Remark 2.2 (Critical Pairs).
The confluence and termination of $\mathrm{LND}_{\mathrm{EQ}}$-TRS are established in [12] via analysis of 56 critical pairs. The key insight is that the canon rule—which rewrites any path $p$ to $\mathrm{ofEq}(p.\mathrm{toEq})$—acts as a “normalization shortcut” that collapses all parallel paths to a common canonical form.
In Lean, the Step relation is defined inductively:
1inductive Step : {A : Type u} -> {a b : A} -> Path a b -> Path a b -> Prop
2 | symm_refl (a : A) : Step (symm (Path.refl a)) (Path.refl a)
3 | symm_symm (p : Path a b) : Step (symm (symm p)) p
4 | trans_refl_left (p : Path a b) : Step (trans (Path.refl a) p) p
5 | trans_refl_right (p : Path a b) : Step (trans p (Path.refl b)) p
6 | trans_symm (p : Path a b) : Step (trans p (symm p)) (Path.refl a)
7 | symm_trans (p : Path a b) : Step (trans (symm p) p) (Path.refl b)
8 | trans_assoc (p : Path a b) (q : Path b c) (r : Path c d) :
9 Step (trans (trans p q) r) (trans p (trans q r))
10 | canon (p : Path a b) : Step p (Path.ofEq p.toEq)
11 -- ... plus type-specific and context rules (50+ total)
2.3. Rewrite Equality
Two paths $p,q:\mathrm{Path}(a,b)$ are rewrite equal ($p\sim q$) if they can be transformed into each other via the $\mathrm{LND}_{\mathrm{EQ}}$-TRS rewrite system. The key rewrite rules include:
$\rho\cdot p$ $\rightsquigarrow p$ (left unit)
$p\cdot\rho$ $\rightsquigarrow p$ (right unit)
$p\cdot p^{-1}$ $\rightsquigarrow\rho$ (right inverse)
$p^{-1}\cdot p$ $\rightsquigarrow\rho$ (left inverse)
$(p\cdot q)\cdot r$ $\rightsquigarrow p\cdot(q\cdot r)$ (associativity)
Example 2.3 (Rewrite Derivation).
Consider $p\cdot(q\cdot q^{-1})$ where $p,q:\mathrm{Path}(a,a)$. The derivation to $p$ proceeds:
$p\cdot(q\cdot q^{-1})$ $\rightsquigarrow p\cdot\rho$ (right inverse)
$\rightsquigarrow p$ (right unit)
This explicit derivation is the “computational content” witnessing that $p\cdot(q\cdot q^{-1})\sim p$.
In Lean, rewrite equality is defined as the equivalence closure of Step:
1inductive RwEq {A : Type u} {a b : A} : Path a b -> Path a b -> Prop
2 | refl (p : Path a b) : RwEq p p
3 | step {p q : Path a b} : Step p q -> RwEq p q
4 | symm {p q : Path a b} : RwEq p q -> RwEq q p
5 | trans {p q r : Path a b} : RwEq p q -> RwEq q r -> RwEq p r
The $\mathrm{LND}_{\mathrm{EQ}}$-TRS enjoys termination and confluence [14, 12], so every path has a unique normal form. This yields three important properties:
  1. (1)
    Soundness: $p\sim q$ implies $p.\mathrm{toEq}=q.\mathrm{toEq}$ (rewriting preserves the underlying equality).
  2. (2)
    Completeness: Any path $p$ is rewrite-equivalent to its canonical form, so $p\sim\mathrm{ofEq}(p.\mathrm{toEq})$.
  3. (3)
    Decidability: $\sim$ is decidable by comparing canonical forms: $p\sim q$ iff their canonical forms are identical.
Theorem 2.4 (Canonicity).
In Lean 4 (with proof-irrelevant Prop), for any paths $p,q:\mathrm{Path}(a,b)$, we have $p\sim q$.
Proof.
Since $\mathrm{Path}(a,b)$ stores a proof $p.\mathrm{toEq}:a=b$ in Lean’s proof-irrelevant $\mathsf{Prop}$, we have $p.\mathrm{toEq}=q.\mathrm{toEq}$ definitionally. The canon rule gives $p\sim\mathrm{ofEq}(p.\mathrm{toEq})$ and $q\sim\mathrm{ofEq}(q.\mathrm{toEq})$. Since $p.\mathrm{toEq}=q.\mathrm{toEq}$, these canonical forms are identical, so $p\sim q$ by transitivity. ∎
Remark 2.5 ($\sim$ Triviality in Lean).
In our Lean 4 implementation, since Path.toEq stores a proof in the proof-irrelevant Prop type, any two paths $p,q:\mathrm{Path}(a,b)$ satisfy $p.\mathrm{toEq}=q.\mathrm{toEq}$ definitionally. Thus $\sim$ is trivially satisfied for all parallel paths. The groupoid laws become derivable rather than requiring explicit rewrite sequences.
This differs from the original computational paths theory where paths carry explicit rewrite traces. The advantage is simpler formalization; the tradeoff is that we lose the ability to distinguish paths by their internal structure. Future work could implement paths as syntactic terms (not carrying equality proofs) to recover the intended distinction.
2.4. Loop Spaces and Fundamental Groups
Definition 2.6 (Path-Connected).
A type $A$ is path-connected if for all $a,b:A$, there exists $p:\mathrm{Path}(a,b)$. Equivalently, $\prod_{a,b:A}\|\mathrm{Path}(a,b)\|_{-1}$ where $\|-\|_{-1}$ denotes propositional truncation.
Definition 2.7 (Loop Space).
The loop space of a type $A$ at a basepoint $a:A$ is:
$\Omega(A,a):=\mathrm{Path}(a,a)$
Definition 2.8 (Fundamental Group).
The fundamental group $\pi_{1}(A,a)$ is the quotient of the loop space by rewrite equality:
$\pi_{1}(A,a):=\Omega(A,a)/{\sim}$
In Lean, the fundamental group is realized as a quotient:
1abbrev LoopSpace (A : Type u) (a : A) : Type u := Path a a
2
3def PiOne (A : Type u) (a : A) : Type u := Quot (@RwEq A a a)
4
5notation "pi_1(" A "," a ")" => PiOne A a
The group operations on $\pi_{1}(A,a)$ are induced from path operations:
  • Identity: $\rho]$
  • Multiplication: $\alpha]\cdot[\beta]:=[\alpha\cdot\beta]$
  • Inverse: $\alpha]^{-1}:=[\alpha^{-1}]$
The group laws (associativity, unit, inverse) hold because the corresponding path identities are witnessed by $\sim$.
Remark 2.9 (Base Point Dependence).
The fundamental group $\pi_{1}(A,a)$ depends on the choice of basepoint $a$. For different basepoints $a,b:A$ in a path-connected type, the fundamental groups are isomorphic via conjugation by a path $\gamma:\mathrm{Path}(a,b)$:
$\pi_{1}(A,a)\cong\pi_{1}(A,b),\quad[\alpha]\mapsto[\gamma^{-1}\cdot\alpha\cdot\gamma]$
However, this isomorphism is not canonical—it depends on the choice of $\gamma$.
3. Pushouts as Higher-Inductive Types
3.1. The Pushout Type
Given types $A$, $B$, $C$ with maps $f:C\to A$ and $g:C\to B$, the pushout $\mathrm{Pushout}(A,B,C,f,g)$ is the higher-inductive type that “glues” $A$ and $B$ together along the common image of $C$:
[…]
Definition 3.1 (Pushout).
The pushout $\mathrm{Pushout}(A,B,C,f,g)$ is characterized by:
  1. (1)
    Point constructors:
    $\mathrm{inl}$ $:A\to\mathrm{Pushout}(A,B,C,f,g)$
    $\mathrm{inr}$ $:B\to\mathrm{Pushout}(A,B,C,f,g)$
  2. (2)
    Path constructor: For each $c:C$,
    $\mathrm{glue}(c):\mathrm{Path}(\mathrm{inl}(f(c)),\mathrm{inr}(g(c)))$
In Lean, pushouts are axiomatized as a HIT:
1axiom Pushout (A : Type u) (B : Type u) (C : Type u)
2 (f : C -> A) (g : C -> B) : Type u
3
4namespace Pushout
5 axiom inl : A -> Pushout A B C f g
6 axiom inr : B -> Pushout A B C f g
7 axiom glue (c : C) : Path (inl (f c)) (inr (g c))
8end Pushout
3.2. Recursion Principle
Definition 3.2 (Pushout Recursion).
Given a type $D$ with:
  • $\mathrm{onInl}:A\to D$
  • $\mathrm{onInr}:B\to D$
  • $\mathrm{onGlue}:\prod_{c:C}\mathrm{Path}(\mathrm{onInl}(f(c)),\mathrm{onInr}(g(c)))$
There exists a function $\mathrm{rec}:\mathrm{Pushout}(A,B,C,f,g)\to D$ such that:
$\mathrm{rec}(\mathrm{inl}(a))$ $=\mathrm{onInl}(a)$
$\mathrm{rec}(\mathrm{inr}(b))$ $=\mathrm{onInr}(b)$
$(\mathrm{rec})_{*}(\mathrm{glue}(c))$ $\sim\mathrm{onGlue}(c)$
3.3. Glue Naturality
A key property for the SVK theorem is the naturality of glue paths:
Lemma 3.3 (Glue Naturality).
For any path $p:\mathrm{Path}(c_{1},c_{2})$ in $C$, the following square commutes up to $\sim$:
$\mathrm{inl}(f(c_{1}))$$\mathrm{inl}(f(c_{2}))$$\mathrm{inr}(g(c_{1}))$$\mathrm{inr}(g(c_{2}))$$\scriptstyle{\mathrm{inl}_{*}(f_{*}(p))}$$\scriptstyle{\mathrm{glue}(c_{1})}$$\scriptstyle{\mathrm{glue}(c_{2})}$$\scriptstyle{\mathrm{inr}_{*}(g_{*}(p))}$
That is, we have the rewrite equality:
$\mathrm{inl}_{*}(f_{*}(p))\cdot\mathrm{glue}(c_{2})\sim\mathrm{glue}(c_{1})\cdot\mathrm{inr}_{*}(g_{*}(p))$
Rearranging, we obtain a form useful for the decode function:
Corollary 3.4.
For any path $p:\mathrm{Path}(c_{1},c_{2})$ in $C$:
$\mathrm{inl}_{*}(f_{*}(p))\sim\mathrm{glue}(c_{1})\cdot\mathrm{inr}_{*}(g_{*}(p))\cdot\mathrm{glue}(c_{2})^{-1}$
In Lean:
1axiom glue_natural {c1 c2 : C} (p : Path c1 c2) :
2 Path.trans (inlPath (Path.congrArg f p)) (glue c2) =
3 Path.trans (glue c1) (inrPath (Path.congrArg g p))
4
5axiom glue_natural_rweq {c1 c2 : C} (p : Path c1 c2) :
6 RwEq (inlPath (Path.congrArg f p))
7 (Path.trans (glue c1)
8 (Path.trans (inrPath (Path.congrArg g p))
9 (Path.symm (glue c2))))
Remark 3.5 (Two Forms of Glue Naturality).
The axiom glue_natural provides a propositional equality of paths, while glue_natural_rweq provides the $\sim$ equivalence needed for quotient arguments. In a cubical type theory, both would be provable; here, we axiomatize them.
3.4. Special Cases
3.4.1. Wedge Sum
The wedge sum $A\vee B$ is the pushout of $A\leftarrow\mathrm{pt}\to B$:
$A\vee B:=\mathrm{Pushout}(A,B,\mathbf{1},\lambda\_.a_{0},\lambda\_.b_{0})$
This identifies the basepoints $a_{0}:A$ and $b_{0}:B$.
Example 3.6 (Figure-Eight as Wedge).
The figure-eight space is $S^{1}\vee S^{1}$—two circles joined at a single point. The glue path identifies the basepoints of the two circles:
$\mathrm{glue}(*):\mathrm{Path}(\mathrm{inl}(\mathrm{base}_{1}),\mathrm{inr}(\mathrm{base}_{2}))$
In Lean:
1def Wedge (A : Type u) (B : Type u) (a0 : A) (b0 : B) : Type u :=
2 Pushout A B PUnit (fun _ => a0) (fun _ => b0)
3.4.2. Suspension
The suspension $\Sigma A$ is the pushout of $\mathbf{1}\leftarrow A\to\mathbf{1}$:
$\Sigma A:=\mathrm{Pushout}(\mathbf{1},\mathbf{1},A,\lambda\_.*,\lambda\_.*)$
This adds a “north pole”, “south pole”, and meridian paths connecting them.
In Lean:
1def Suspension (A : Type u) : Type u :=
2 Pushout PUnit PUnit A (fun _ => PUnit’.unit) (fun _ => PUnit’.unit)
3
4noncomputable def north : Suspension A := Pushout.inl PUnit’.unit
5noncomputable def south : Suspension A := Pushout.inr PUnit’.unit
6noncomputable def merid (a : A) : Path north south := Pushout.glue a
4. Free Products and Amalgamated Free Products
4.1. Free Product Words
Definition 4.1 (Free Product Word).
A word in the free product $G_{1}\ast G_{2}$ is an alternating sequence of elements from $G_{1}$ and $G_{2}$:
1inductive FreeProductWord (G1 : Type u) (G2 : Type u) : Type u
2 | nil : FreeProductWord G1 G2
3 | consLeft (x : G1) (rest : FreeProductWord G1 G2) : FreeProductWord G1 G2
4 | consRight (y : G2) (rest : FreeProductWord G1 G2) : FreeProductWord G1 G2
Example 4.2 (Figure-Eight Words).
For $\pi_{1}(S^{1}\vee S^{1})\simeq\mathbb{Z}\ast\mathbb{Z}$, words represent sequences of loops around the two circles:
  • $\mathrm{consLeft}(1,\mathrm{nil})$ — one loop around the left circle
  • $\mathrm{consRight}(2,\mathrm{nil})$ — two loops around the right circle
  • $\mathrm{consLeft}(1,\mathrm{consRight}(-1,\mathrm{nil}))$ — left loop, then reverse right loop
Words support concatenation:
1def concat : FreeProductWord G1 G2 -> FreeProductWord G1 G2
2 -> FreeProductWord G1 G2
3 | nil, w2 => w2
4 | consLeft x rest, w2 => consLeft x (concat rest w2)
5 | consRight y rest, w2 => consRight y (concat rest w2)
Lemma 4.3 (Word Concatenation Properties).
  1. (1)
    $\mathrm{concat}(\mathrm{nil},w)=w$
  2. (2)
    $\mathrm{concat}(w,\mathrm{nil})=w$
  3. (3)
    $\mathrm{concat}(\mathrm{concat}(w_{1},w_{2}),w_{3})=\mathrm{concat}(w_{1},\mathrm{concat}(w_{2},w_{3}))$
Free vs. Direct Product.
Unlike the torus $S^{1}\times S^{1}$ whose loops can “slide past each other” (yielding the abelian group $\mathbb{Z}\times\mathbb{Z}$), loops in $S^{1}\vee S^{1}$ must pass through the common basepoint, preventing commutation. The word $\mathrm{loop}_{A}\cdot\mathrm{loop}_{B}$ represents first traversing the left circle, then the right; reversing the order gives a genuinely distinct homotopy class $\mathrm{loop}_{B}\cdot\mathrm{loop}_{A}$.
4.2. Amalgamated Free Product
When $G_{1}$ and $G_{2}$ share a common subgroup $H$ via homomorphisms $i_{1}:H\to G_{1}$ and $i_{2}:H\to G_{2}$, the amalgamated free product $G_{1}\ast_{H}G_{2}$ identifies $i_{1}(h)$ with $i_{2}(h)$ for all $h:H$.
Definition 4.4 (Amalgamation Relation).
The amalgamation relation identifies elements from $H$ mapped into $G_{1}$ with their counterparts mapped into $G_{2}$:
1inductive AmalgRelation {G1 G2 H : Type u}
2 (i1 : H -> G1) (i2 : H -> G2) :
3 FreeProductWord G1 G2 -> FreeProductWord G1 G2 -> Prop
4 | amalgLeftToRight (h : H) (pre suf : FreeProductWord G1 G2) :
5 AmalgRelation i1 i2
6 (concat pre (concat (singleLeft (i1 h)) suf))
7 (concat pre (concat (singleRight (i2 h)) suf))
8 | amalgRightToLeft (h : H) (pre suf : FreeProductWord G1 G2) :
9 AmalgRelation i1 i2
10 (concat pre (concat (singleRight (i2 h)) suf))
11 (concat pre (concat (singleLeft (i1 h)) suf))
Definition 4.5 (Amalgamated Free Product).
The amalgamated free product is the quotient of free product words by the equivalence closure of the amalgamation relation:
1def AmalgamatedFreeProduct (G1 G2 H : Type u)
2 (i1 : H -> G1) (i2 : H -> G2) : Type u :=
3 Quot (AmalgEquiv i1 i2)
where $\mathrm{AmalgEquiv}$ is the equivalence closure of $\mathrm{AmalgRelation}$.
Remark 4.6 (Structure of AmalgEquiv).
The relation $\mathrm{AmalgEquiv}$ includes:
  1. (1)
    Standard word-reduction equivalences for the free product $G_{1}\ast G_{2}$: combining adjacent letters from the same factor, and removing identity elements.
  2. (2)
    The amalgamation relations identifying $i_{1}(h)$ with $i_{2}(h)$ for all $h:H$.
  3. (3)
    Reflexive, symmetric, and transitive closure.
In the Lean formalization, we build the free product quotient first, then extend with amalgamation. The result is the pushout in the category of groups.
Lemma 4.7 (Amalgamation Property).
For any $h:H$:
$\mathrm{inl}(i_{1}(h))=\mathrm{inr}(i_{2}(h))\quad\text{in }G_{1}\ast_{H}G_{2}$
Proof.
By definition, the singleton words $\mathrm{singleLeft}(i_{1}(h))$ and $\mathrm{singleRight}(i_{2}(h))$ are related by $\mathrm{AmalgRelation}$ (with empty prefix and suffix). The quotient $\mathrm{Quot.sound}$ then yields the equality. ∎
5. The Seifert-van Kampen Theorem
With pushouts (Section 3) and amalgamated free products (Section 4) in hand, we can now state and prove the main theorem.
5.1. Statement
Theorem 5.1 (Seifert-van Kampen).
Let $A$, $B$, $C$ be path-connected types with maps $f:C\to A$ and $g:C\to B$, and let $c_{0}:C$. Then:
$\pi_{1}(\mathrm{Pushout}(A,B,C,f,g),\mathrm{inl}(f(c_{0})))\simeq\pi_{1}(A,f(c_{0}))\ast_{\pi_{1}(C,c_{0})}\pi_{1}(B,g(c_{0}))$
where the amalgamation is via the induced maps $f_{*}:\pi_{1}(C,c_{0})\to\pi_{1}(A,f(c_{0}))$ and $g_{*}:\pi_{1}(C,c_{0})\to\pi_{1}(B,g(c_{0}))$.
Proof sketch.
We construct:
  1. (1)
    $\mathrm{decode}$: Maps words to loops by interpreting left letters as loops in $\mathrm{inl}(A)$, and right letters as loops in $\mathrm{inr}(B)$ conjugated by $\mathrm{glue}(c_{0})$.
  2. (2)
    $\mathrm{encode}$: Defined via transport of a code family along paths (axiomatized).
The key lemmas are:
  • Decode respects amalgamation (Lemma 5.3): elements from $\pi_{1}(C)$ decode identically whether placed in the left or right factor.
  • Round-trip properties: $\mathrm{decode}\circ\mathrm{encode}=\mathrm{id}$ and $\mathrm{encode}\circ\mathrm{decode}\sim_{\mathrm{amalg}}\mathrm{id}$.
The first round-trip follows by path induction on pushout paths; the second by word induction. See §5.2.35.2.4 for details. ∎
5.2. The Encode-Decode Method
The proof uses the encode-decode method:
  1. (1)
    Decode: Define a map from the amalgamated free product to $\pi_{1}(\mathrm{Pushout})$.
  2. (2)
    Encode: Define a map from $\pi_{1}(\mathrm{Pushout})$ to the amalgamated free product.
  3. (3)
    Round-trip: Show both compositions are identity.
5.2.1. The Decode Map
Definition 5.2 (Pushout Decode).
The decode map converts a word to a loop in the pushout:
1noncomputable def pushoutDecode (c0 : C) :
2 FreeProductWord (pi_1(A, f c0)) (pi_1(B, g c0))
3 -> pi_1(Pushout A B C f g, Pushout.inl (f c0))
4 | .nil => Quot.mk _ (Path.refl _)
5 | .consLeft alpha rest =>
6 let alphaInPushout := Quot.lift
7 (fun p => Quot.mk RwEq (Pushout.inlPath p))
8 (fun _ _ hp => Quot.sound (rweq_congrArg_of_rweq Pushout.inl hp))
9 alpha
10 piOneMul alphaInPushout (pushoutDecode c0 rest)
11 | .consRight beta rest =>
12 let betaInPushout := Quot.lift
13 (fun q => Quot.mk RwEq (Path.trans
14 (Pushout.glue c0)
15 (Path.trans (Pushout.inrPath q)
16 (Path.symm (Pushout.glue c0)))))
17 (fun _ _ hq => Quot.sound (...))
18 beta
19 piOneMul betaInPushout (pushoutDecode c0 rest)
The key insight for elements from $\pi_{1}(B)$: since the basepoint is in $\mathrm{inl}(A)$, we must conjugate by the glue path to create a loop:
$\beta\mapsto\mathrm{glue}(c_{0})\cdot\mathrm{inr}_{*}(\beta)\cdot\mathrm{glue}(c_{0})^{-1}$
$\mathrm{inl}(f(c_{0}))$$\mathrm{inr}(g(c_{0}))$$\mathrm{inl}(f(c_{0}))$$\mathrm{inr}(g(c_{0}))$$\scriptstyle{\mathrm{glue}(c_{0})}$$\scriptstyle{\beta_{\mathrm{decoded}}}$$\scriptstyle{\mathrm{inr}_{*}(\beta)}$$\scriptstyle{\mathrm{glue}(c_{0})^{-1}}$
5.2.2. Decode Respects Amalgamation
Lemma 5.3 (Decode Respects Amalgamation).
For any $\gamma:\pi_{1}(C,c_{0})$ and word $w$:
$\mathrm{decode}(\mathrm{consLeft}(f_{*}(\gamma),w))=\mathrm{decode}(\mathrm{consRight}(g_{*}(\gamma),w))$
Proof.
We need to show that for a loop $p$ in $C$ at $c_{0}$:
$\mathrm{inl}_{*}(f_{*}(p))\sim\mathrm{glue}(c_{0})\cdot\mathrm{inr}_{*}(g_{*}(p))\cdot\mathrm{glue}(c_{0})^{-1}$
This is exactly Corollary 3.4 applied to the loop $p$ (where $c_{1}=c_{2}=c_{0}$):
1theorem glue_natural_loop_rweq (c0 : C) (p : LoopSpace C c0) :
2 RwEq (inlPath (Path.congrArg f p))
3 (Path.trans (glue c0)
4 (Path.trans (inrPath (Path.congrArg g p))
5 (Path.symm (glue c0)))) :=
6 glue_natural_rweq p
5.2.3. The Encode Map
The encode map uses the “code family” approach, a technique central to HoTT calculations of fundamental groups [10, 17]. The idea is to define a type family $\mathrm{Code}:\mathrm{Pushout}\to\mathrm{Type}$ that, at each point, represents the words that “should decode to a path ending at that point.”
  • At $\mathrm{inl}(a)$: $\mathrm{Code}(\mathrm{inl}(a))$ represents words decoding to paths ending at $\mathrm{inl}(a)$
  • At $\mathrm{inr}(b)$: $\mathrm{Code}(\mathrm{inr}(b))$ represents words decoding to paths ending at $\mathrm{inr}(b)$
  • Transport along $\mathrm{glue}(c)$: adjusts words by prepending/removing letters for the glue path
The encode function is then:
$\mathrm{encode}(p):=\mathrm{transport}(\mathrm{Code},p,\mathrm{nil})$
Why encode is axiomatized.
Fully constructing the code family requires dependent elimination over the pushout HIT, which in Lean 4 (without native HIT support or cubical features) would require extensive additional infrastructure. Since our focus is on the computational paths framework, we axiomatize encode’s existence and key properties:
1axiom pushoutEncodeAxiom (A B C : Type u) (f : C -> A) (g : C -> B) (c0 : C) :
2 LoopSpace (Pushout A B C f g) (Pushout.inl (f c0)) ->
3 FreeProductWord (pi_1(A, f c0)) (pi_1(B, g c0))
4
5axiom pushoutEncodeAxiom_respects_rweq (A B C : Type u)
6 (f : C -> A) (g : C -> B) (c0 : C)
7 {p q : LoopSpace (Pushout A B C f g) (Pushout.inl (f c0))}
8 (h : RwEq p q) :
9 pushoutEncodeAxiom A B C f g c0 p = pushoutEncodeAxiom A B C f g c0 q
The second axiom ensures encode respects $\sim$, essential for lifting to the $\pi_{1}$ quotient. In cubical type theory, these would be provable; here, they capture the interface needed for SVK.
5.2.4. Round-Trip Properties
Lemma 5.4 (Decode-Encode Round-Trip).
For any loop $p$ in the pushout at $\mathrm{inl}(f(c_{0}))$:
$\mathrm{decode}(\mathrm{encode}(p))=[p]$
Lemma 5.5 (Encode-Decode Round-Trip).
For any word $w$:
$\mathrm{encode}(\mathrm{decode}(w))\sim_{\mathrm{amalg}}w$
where $\sim_{\mathrm{amalg}}$ is amalgamation equivalence.
5.3. The Equivalence
The equivalence is defined in Lean as:
1noncomputable def seifertVanKampenEquiv :
2 SimpleEquiv
3 (pi_1(Pushout A B C f g, Pushout.inl (f c0)))
4 (AmalgamatedFreeProduct
5 (pi_1(A, f c0))
6 (pi_1(B, g c0))
7 (pi_1(C, c0))
8 (piOneFmap c0)
9 (piOneGmap c0)) where
10 toFun := pushoutEncodeAmalg c0
11 invFun := pushoutDecodeAmalg c0
12 left_inv := by
13 intro alpha
14 induction alpha using Quot.ind with
15 | _ p => exact pushoutDecodeEncodeAxiom A B C f g c0 p
16 right_inv := by
17 intro x
18 induction x using Quot.ind with
19 | _ w =>
20 apply Quot.sound
21 exact pushoutEncodeDecodeAxiom A B C f g c0 w
6. Applications
6.1. The Wedge Sum: Free Product of Fundamental Groups
When the gluing space $C$ is a point, $\pi_{1}(C)$ is trivial and the amalgamated free product reduces to a true free product.
Theorem 6.1 (Fundamental Group of Wedge Sum).
For pointed types $(A,a_{0})$ and $(B,b_{0})$:
$\pi_{1}(A\vee B)\simeq\pi_{1}(A)\ast\pi_{1}(B)$
Proof.
The wedge $A\vee B=\mathrm{Pushout}(A,B,\mathbf{1},\lambda\_.a_{0},\lambda\_.b_{0})$. By SVK:
$\pi_{1}(A\vee B)\simeq\pi_{1}(A)\ast_{\pi_{1}(\mathbf{1})}\pi_{1}(B)$
Since $\pi_{1}(\mathbf{1})\cong 1$ (the trivial group), the amalgamation imposes no relations, yielding the free product. ∎
In Lean:
1noncomputable def wedgeFundamentalGroupEquiv :
2 SimpleEquiv
3 (pi_1(Wedge A B a0 b0, Wedge.basepoint))
4 (FreeProductWord (pi_1(A, a0)) (pi_1(B, b0))) where
5 toFun := wedgeEncodeQuot a0 b0
6 invFun := wedgeFreeProductDecode a0 b0
7 left_inv := ...
8 right_inv := ...
6.2. The Figure-Eight Space
Theorem 6.2 (Fundamental Group of Figure-Eight).
$\pi_{1}(S^{1}\vee S^{1})\simeq\mathbb{Z}\ast\mathbb{Z}$
Proof.
The figure-eight is the wedge of two circles: $\mathrm{FigureEight}=S^{1}\vee S^{1}$.
By Theorem 6.1:
$\pi_{1}(S^{1}\vee S^{1})\simeq\pi_{1}(S^{1})\ast\pi_{1}(S^{1})\simeq\mathbb{Z}\ast\mathbb{Z}$
using the fundamental theorem $\pi_{1}(S^{1})\simeq\mathbb{Z}$ [10]. ∎
In Lean:
1def FigureEight : Type u := Wedge Circle Circle circleBase circleBase
2
3noncomputable def figureEightPiOneEquivPiOneProduct :
4 SimpleEquiv FigureEightPiOne (FreeProductWord circlePiOne circlePiOne) :=
5 wedgeFundamentalGroupEquiv circleBase circleBase
6
7noncomputable def figureEightPiOneEquiv :
8 SimpleEquiv FigureEightPiOne (FreeProductWord Int Int) :=
9 SimpleEquiv.comp figureEightPiOneEquivPiOneProduct piOneProductEquivIntProduct
6.2.1. Non-Abelianness
The figure-eight has a non-abelian fundamental group, witnessed by the word structure:
1def wordAB : FreeProductWord Int Int := .consLeft 1 (.consRight 1 .nil)
2def wordBA : FreeProductWord Int Int := .consRight 1 (.consLeft 1 .nil)
3
4theorem wordAB_ne_wordBA : wordAB != wordBA := by
5 intro h
6 cases h -- consLeft and consRight are distinct constructors
This shows that $\mathrm{loop}_{A}\cdot\mathrm{loop}_{B}\neq\mathrm{loop}_{B}\cdot\mathrm{loop}_{A}$. The proof is by case analysis: FreeProductWord is an inductive type where consLeft and consRight are distinct constructors, so no sequence of reductions can make them equal.
Topologically, $\mathrm{wordAB}$ corresponds to first traversing the left circle, then the right; $\mathrm{wordBA}$ traverses them in the opposite order. In the figure-eight, these loops cannot be continuously deformed into each other—they are genuinely distinct homotopy classes.
6.3. The 2-Sphere: Trivial Fundamental Group
Theorem 6.3 (Fundamental Group of 2-Sphere).
$\pi_{1}(S^{2})\simeq 1$
Proof.
The 2-sphere is the suspension of the circle: $S^{2}=\Sigma(S^{1})$.
As a pushout: $\Sigma(S^{1})=\mathrm{Pushout}(\mathbf{1},\mathbf{1},S^{1},!,!)$ where !! is the unique map to the unit type.
By SVK:
$\pi_{1}(S^{2})\simeq\pi_{1}(\mathbf{1})\ast_{\pi_{1}(S^{1})}\pi_{1}(\mathbf{1})=1\ast_{\mathbb{Z}}1=1$
The key insight: when both factors are trivial, the amalgamated product is trivial regardless of the amalgamating group. ∎
Lemma 6.4 (Trivial Decode).
When $\pi_{1}(A)=\pi_{1}(B)=1$, for any word $w$:
$\mathrm{decode}(w)=[\rho]$
Proof.
Each letter in $w$ comes from either $\pi_{1}(A)$ or $\pi_{1}(B)$, both trivial. Thus each letter represents the identity loop $\rho]$, and $\mathrm{decode}(w)=[\rho]^{n}=[\rho]$. ∎
In Lean:
1theorem trivial_decode (f : C -> PUnit’) (g : C -> PUnit’) (c0 : C)
2 (w : FreeProductWord (pi_1(PUnit’, PUnit’.unit)) (pi_1(PUnit’, PUnit’.unit))) :
3 pushoutDecode c0 w = Quot.mk _ (Path.refl _) := by
4 induction w with
5 | nil => rfl
6 | consLeft alpha rest ih =>
7 simp only [pushoutDecode]
8 rw [trivial_left_inclusion c0 alpha]
9 rw [ih]
10 exact piOneMul_refl_left _
11 | consRight beta rest ih =>
12 simp only [pushoutDecode]
13 rw [trivial_right_inclusion c0 beta]
14 rw [ih]
15 exact piOneMul_refl_left _
16
17theorem sphere2_pi1_trivial :
18 forall (alpha : pi_1(Sphere2, basepoint)), alpha = Quot.mk _ (Path.refl _) := ...
19
20noncomputable def sphere2_pi1_equiv_unit :
21 SimpleEquiv (pi_1(Sphere2, basepoint)) Unit where
22 toFun := fun _ => ()
23 invFun := fun _ => Quot.mk _ (Path.refl _)
24 left_inv := by intro alpha; exact (sphere2_pi1_trivial alpha).symm
25 right_inv := by intro u; cases u; rfl
Remark 6.5 (Higher Homotopy Groups).
The SVK theorem has no direct analogue for higher homotopy groups: $\pi_{n}(A\vee B)\not\simeq\pi_{n}(A)\times\pi_{n}(B)$ in general for $n>1$. Instead, the Blakers-Massey theorem [8] provides connectivity bounds. Computing higher $\pi_{n}$ via computational paths would require the weak $\omega$-groupoid structure established in prior work [5].
6.4. The 3-Sphere and the Hopf Fibration
The SVK technique for proving $\pi_{1}(S^{2})\simeq 1$ generalizes immediately to higher-dimensional spheres. We demonstrate this with the 3-sphere.
Theorem 6.6 (Fundamental Group of 3-Sphere).
$\pi_{1}(S^{3})\simeq 1$
Proof.
The 3-sphere is the suspension of the 2-sphere: $S^{3}=\Sigma(S^{2})$.
As a pushout: $\Sigma(S^{2})=\mathrm{Pushout}(\mathbf{1},\mathbf{1},S^{2},!,!)$.
By SVK:
$\pi_{1}(S^{3})\simeq\pi_{1}(\mathbf{1})\ast_{\pi_{1}(S^{2})}\pi_{1}(\mathbf{1})=1\ast_{1}1=1$
The proof is identical in structure to Theorem 6.3—only the amalgamating group changes from $\pi_{1}(S^{1})\cong\mathbb{Z}$ to $\pi_{1}(S^{2})\cong 1$. Since both factors $\pi_{1}(\mathbf{1})$ are trivial, the result is trivial regardless. ∎
This pattern continues: for any $n\geq 2$, we have $S^{n}=\Sigma(S^{n-1})$, and by induction $\pi_{1}(S^{n})\simeq 1$.
In Lean, the proof reuses the same SVK machinery:
1theorem sphere3_pi1_trivial :
2 forall (l : LoopSpace Sphere3 sphere3North),
3 Quot.mk RwEq l = Quot.mk RwEq (Path.refl sphere3North) := by
4 intro l
5 let f : Sphere2 -> PUnit := sphere2ToNorth
6 let g : Sphere2 -> PUnit := sphere2ToSouth
7 let c0 : Sphere2 := basepoint
8 let encoded := pushoutEncodeAmalg c0 (Quot.mk RwEq l)
9 have left_inv_l : Quot.mk RwEq l = pushoutDecodeAmalg c0 encoded := by
10 simp only [pushoutDecodeAmalg]
11 exact (pushoutDecodeEncodeAxiom PUnit PUnit Sphere2 f g c0 l).symm
12 rw [left_inv_l]
13 exact amalg_trivial_is_one_s3 encoded
6.4.1. The Hopf Fibration Context
The 3-sphere is the total space of the Hopf fibration, a fundamental fiber bundle in topology:
$S^{1}\longrightarrow S^{3}\longrightarrow S^{2}$
This bundle has fiber $S^{1}$, total space $S^{3}$, and base $S^{2}$.
In the computational paths framework, we model this via a type family:
  • Hopf fiber family: $P:S^{2}\to\mathcal{U}$ with $P(x)=S^{1}$ for all $x$
  • Total space: $\Sigma_{x:S^{2}}P(x)$, equivalent to $S^{3}$
  • Hopf projection: $\mathrm{hopfProj}:\Sigma_{x:S^{2}}P(x)\to S^{2}$
The fiber equivalence $\mathrm{Fiber}(\mathrm{hopfProj},x)\simeq S^{1}$ is proved (not axiomatized) from the $\Sigma$-type construction.
Remark 6.7 (Long Exact Sequence).
The Hopf fibration induces a long exact sequence of homotopy groups:
$\cdots\to\pi_{2}(S^{1})\to\pi_{2}(S^{3})\to\pi_{2}(S^{2})\xrightarrow{\partial}\pi_{1}(S^{1})\to\pi_{1}(S^{3})\to\pi_{1}(S^{2})\to\cdots$
Substituting known values:
$\cdots\to 1\to\pi_{2}(S^{3})\to\pi_{2}(S^{2})\xrightarrow{\partial}\mathbb{Z}\to 1\to 1$
By exactness, the connecting map $\partial:\pi_{2}(S^{2})\to\mathbb{Z}$ is an isomorphism (in untruncated homotopy theory), yielding $\pi_{2}(S^{2})\cong\mathbb{Z}$. In our truncated framework where $\pi_{2}\cong 1$, this collapses, but the fibration structure remains valuable for understanding the topology.
Remark 6.8 (Non-Triviality).
The Hopf fibration is non-trivial: $S^{3}\not\simeq S^{2}\times S^{1}$. If it were a trivial product, we would have $\pi_{1}(S^{3})\cong\pi_{1}(S^{2})\times\pi_{1}(S^{1})\cong 1\times\mathbb{Z}\cong\mathbb{Z}$, contradicting $\pi_{1}(S^{3})\cong 1$.
7. The Lean 4 Formalization
7.1. Architecture
The Lean 4 formalization is organized as follows:
Module Content
Path/Basic/Core.lean Path type, refl, symm, trans
Path/Rewrite/Step.lean Single-step rewrite rules (50+)
Path/Rewrite/RwEq.lean Rewrite equality relation
Path/Homotopy/FundamentalGroup.lean $\pi_{1}$ definition
Path/HIT/Pushout.lean Pushout HIT axioms
Path/HIT/PushoutPaths.lean SVK theorem
Path/HIT/FigureEight.lean Figure-eight application
Path/HIT/Sphere.lean 2-sphere application
Path/HIT/Torus.lean $\pi_{1}(T^{2})\simeq\mathbb{Z}\times\mathbb{Z}$
Path/HIT/KleinBottle.lean $\pi_{1}(K)\simeq\mathbb{Z}\rtimes\mathbb{Z}$
Path/HIT/OrientableSurface.lean $\pi_{1}(\Sigma_{g})$ for genus $g$
Path/HIT/HopfFibration.lean $S^{3}$, Hopf map, $\pi_{1}(S^{3})=1$
Path/Homotopy/Fibration.lean Fiber sequences, long exact sequence
Path/Homotopy/EilenbergMacLane.lean $K(G,n)$ spaces
7.2. Key Definitions
The formalization uses Lean 4’s quotient types for both $\pi_{1}$ (quotient by $\sim$) and the amalgamated free product (quotient by amalgamation equivalence).
1-- Fundamental group as quotient
2def PiOne (A : Type u) (a : A) : Type u := Quot (@RwEq A a a)
3
4-- Amalgamated free product as quotient
5def AmalgamatedFreeProduct (G1 G2 H : Type u)
6 (i1 : H -> G1) (i2 : H -> G2) : Type u :=
7 Quot (AmalgEquiv i1 i2)
7.3. Axiom Inventory
Table 1. Axiom usage in the formalization
Axiom Purpose Justification
Pushout Pushout type exists Standard HIT
Pushout.inl/inr Point constructors Standard HIT
Pushout.glue Path constructor Standard HIT
glue_natural_rweq Naturality of glue Derivable in cubical TT
pushoutEncodeAxiom Encode function exists Code family construction
pushoutEncodeAxiom_respects_rweq Encode respects $\sim$ From encode construction
pushoutDecodeEncodeAxiom Round-trip property Path induction
pushoutEncodeDecodeAxiom Round-trip property Word induction
Remark 7.1 (Noncomputability).
Definitions involving HITs are marked noncomputable because Lean 4 lacks native computational support for higher-inductive types. The decode function, while marked noncomputable, has constructive content—given a word, it produces an explicit loop term. The encode function is genuinely nonconstructive in our axiomatization.
Remark 7.2 (Consistency).
The pushout axioms are standard HIT axioms known to be consistent with MLTT via various models (simplicial sets, cubical sets [1]). The encode axioms are consequences of the full HIT elimination principle, which is derivable in cubical type theories.
7.4. What Is Proved vs. Axiomatized
Table 2. Proof status of main results
Result Status Dependencies
Decode function Proved None
Decode respects amalgamation Proved glue_natural_rweq
Encode function Axiomatized pushoutEncodeAxiom
Round-trip ($\mathrm{decode}\circ\mathrm{encode}$) Axiomatized pushoutDecodeEncodeAxiom
Round-trip ($\mathrm{encode}\circ\mathrm{decode}$) Axiomatized pushoutEncodeDecodeAxiom
SVK equivalence Proved Above axioms
Wedge sum theorem Proved SVK + triviality
Figure-eight $\pi_{1}$ Proved Wedge + $\pi_{1}(S^{1})\simeq\mathbb{Z}$
2-sphere $\pi_{1}$ Proved SVK + triviality
3-sphere $\pi_{1}$ Proved SVK + triviality (same as $S^{2}$)
Non-abelianness of $\mathbb{Z}*\mathbb{Z}$ Proved Definitional (constructor ineq.)
7.5. Statistics
Component Lines of Lean
Pushout definition and lemmas 365
SVK theorem and infrastructure 1,004
Figure-eight application 330
2-sphere application 398
3-sphere (Hopf fibration) 393
Total (SVK-specific) 2,490
Note: This count covers only the SVK-specific modules. The full computational paths library—including path basics, rewrite system, homotopy theory, additional HITs (circle, torus, Klein bottle, projective plane, orientable surfaces), fibrations, Eilenberg-MacLane spaces, and the Hopf fibration—comprises approximately 25,100 lines.
7.6. Automation Potential
Since $\sim$ is decidable via normalization, a Lean tactic could automatically discharge $\pi_{1}$ equality goals by normalizing both sides and comparing. This would automate much of the tedious case analysis in proofs like Lemma 5.3. Future work could implement such a tactic using Lean 4’s metaprogramming facilities.
8. Conclusion
We have presented a formalization of the Seifert-van Kampen theorem within the computational paths framework, demonstrating that:
  1. (1)
    Pushouts can be defined as higher-inductive types with explicit path constructors.
  2. (2)
    The encode-decode method for computing $\pi_{1}(\mathrm{Pushout})$ works smoothly with computational paths, where path equality is captured by the decidable $\sim$ relation.
  3. (3)
    The SVK theorem yields practical applications: the figure-eight has non-abelian fundamental group $\mathbb{Z}\ast\mathbb{Z}$, and the 2-sphere is simply connected.
  4. (4)
    The entire development can be machine-checked in Lean 4, with axioms clearly identified.
8.1. Comparison with HoTT Approaches
Table 3. Comparison with HoTT Book SVK (§8.7)
Aspect Standard HoTT Computational Paths
Path equality Identity type ($\mathrm{Id}$) Decidable ($\sim$)
Coherence witnesses Abstract existence Explicit rewrite derivations
Encode function Code family via $J$ Axiomatized
Decode function Direct construction Direct construction
$\pi_{1}$ computation Quotient Quotient by $\sim$
8.2. Limitations
  1. (1)
    Axiom usage: The encode function and its properties are axiomatized. A full construction would require dependent elimination for HITs, not natively supported in Lean 4. In cubical Agda, these axioms would be provable.
  2. (2)
    Path-connectedness assumption: The SVK theorem requires path-connected types. For spaces with multiple components, a groupoid-valued version would be needed.
  3. (3)
    Rewrite system complexity: While the $\mathrm{LND}_{\mathrm{EQ}}$-TRS has 50+ rules, normalization complexity has not been formally analyzed.
  4. (4)
    Non-computational encode: Due to axiomatization, encode is not computationally evaluable. Decode, in contrast, is fully constructive.
  5. (5)
    $\sim$ triviality: As noted in Remark 2.5, our Lean implementation collapses $\sim$ via proof irrelevance, losing some intended distinctions between paths.
8.3. Recent Extensions
Since the initial development, the codebase has been extended with several features originally planned as future work:
  1. (1)
    Higher homotopy groups $\pi_{n}$: Building on the weak $\omega$-groupoid structure, we now define:
    • Iterated loop spaces: $\texttt{Loop2Space}=\mathcal{D}_{2}(\rho_{a},\rho_{a})$, etc.
    • The second homotopy group $\pi_{2}(A,a)$ as the quotient of 2-loops by $\mathcal{D}_{3}$-connectivity
    • Eckmann-Hilton argument proving $\pi_{2}$ is abelian (piTwo_comm)
    • $\pi_{2}(S^{2})\cong 1$ via contractibility3
  2. (2)
    Truncation levels (n-types): Connecting the $\omega$-groupoid to HoTT truncation theory:
    • IsContr: contractible types (center + contraction paths)
    • IsProp: propositions (any two elements path-connected)
    • IsSet: sets (parallel paths are $\sim$)
    • IsGroupoid: 1-groupoids (parallel derivations connected)
    • Cumulative hierarchy: contractible $\to$ proposition $\to$ set $\to$ groupoid
    • All types are automatically 1-groupoids via contractibility3
  3. (3)
    Path simplification tactic: The path_simp tactic automates $\sim$ proofs:
    1macro "path_simp" : tactic =>
    2 ‘(tactic| simp only [rweq_refl, rweq_cmpA_refl_left,
    3 rweq_cmpA_refl_right, rweq_cmpA_inv_left, ...])
    Additional tactics include path_rfl, path_canon, and path_decide.
  4. (4)
    Covering space theory: Basic covering space infrastructure:
    • Total space $\Sigma_{a:A}P(a)$ for type families $P:A\to\mathcal{U}$
    • Path lifting via transport
    • $\pi_{1}(A,a)$ action on fibers via fiberAction
    • Universal cover with fiber $\pi_{1}(A,a)$
    • Deck transformations as covering automorphisms
  5. (5)
    Encode-decode improvements: The wedge encode-decode proofs now derive injectivity and surjectivity from the fundamental axioms, providing a cleaner interface.
  6. (6)
    Additional fundamental group computations: Beyond the figure-eight and 2-sphere:
    • Torus: $\pi_{1}(T^{2})\simeq\mathbb{Z}\times\mathbb{Z}$ via encode-decode
    • Klein bottle: $\pi_{1}(K)\simeq\mathbb{Z}\rtimes\mathbb{Z}$ (semidirect product)
    • Real projective plane: $\pi_{1}(\mathbb{RP}^{2})\simeq\mathbb{Z}_{2}$
    • Orientable surfaces: $\pi_{1}(\Sigma_{g})\simeq\langle a_{1},b_{1},\ldots,a_{g},b_{g}\mid\prod[a_{i},b_{i}]=1\rangle$
    • Möbius band and cylinder: $\pi_{1}\simeq\mathbb{Z}$ via deformation retract to circle
  7. (7)
    Fibrations and fiber sequences: Full fibration theory for computational paths:
    • Fiber types: $\mathrm{Fiber}(f,b):=\{a:A\mid f(a)=b\}$ with projection and inclusion
    • Fiber sequences $F\to E\to B$ with equivalence to fibers
    • Type families as fibrations with path lifting via transport
    • Connecting map $\partial:\pi_{1}(B,b)\to F$ via transport
    • Long exact sequence: $\pi_{1}(F)\to\pi_{1}(E)\to\pi_{1}(B)\to\pi_{0}(F)$
    • Induced maps $f_{*}:\pi_{1}(A,a)\to\pi_{1}(B,f(a))$ with functoriality
    • Exactness proofs using $\Sigma$-type path characterization
  8. (8)
    Suspension-loop adjunction: Foundations for the adjunction $\Sigma X,Y]_{*}\cong[X,\Omega Y]_{*}$:
    • Pointed types and pointed maps with basepoint preservation
    • Suspension $\Sigma X$ as pointed type (north as basepoint)
    • Loop space $\Omega Y$ as pointed type (refl as basepoint)
    • Adjunction map: $f:\Sigma X\to Y$ yields $X\to\Omega Y$ via meridians
    • Basepoint preservation: $x_{0}\mapsto\rho$ when $f(\mathrm{north})=y_{0}$
    • Connectivity structures: IsPathConnectedPointed, IsSimplyConnected
  9. (9)
    Eilenberg-MacLane spaces $K(G,n)$: Characterization of classifying spaces:
    • Group and abelian group structures (GroupStr, AbelianGroupStr)
    • $K(G,1)$ characterization via IsKG1: path-connected, $\pi_{1}\cong G$, $\pi_{2}=1$
    • Circle is $K(\mathbb{Z},1)$: uses encode-decode with $\pi_{2}$ triviality from contractibility3
    • Loop space property: $\Omega(K(G,n+1))\simeq K(G,n)$ (stated)
    • Classifying spaces: $BG=K(G,1)$ structure
8.4. Remaining Future Work
  1. (1)
    Cubical type theory comparison: Relate computational paths to cubical approaches [1], exploring whether explicit rewrites complement cubical computation.
  2. (2)
    Fully constructive encode: Implement the code family to eliminate encode axioms, via cubical embedding or macro-generated eliminators.
  3. (3)
    Syntactic paths: Implement paths as syntactic terms (rather than equality-proof wrappers) to recover non-trivial $\sim$ distinctions.
  4. (4)
    Covering space classification: Prove the Galois correspondence between connected covers and subgroups of $\pi_{1}$.
  5. (5)
    Hopf fibration completion: The Hopf fibration framework is established with $\pi_{1}(S^{3})=1$ proved via SVK. Remaining work includes: constructing the Hopf map explicitly via the join $S^{1}*S^{1}\simeq S^{3}$; proving the equivalence $S^{3}\simeq\Sigma_{x:S^{2}}S^{1}$ without axioms; and extending to the quaternionic ($S^{3}\to S^{7}\to S^{4}$) and octonionic ($S^{7}\to S^{15}\to S^{8}$) Hopf fibrations.