### v3

# The Seifert-van Kampen Theorem via Computational Paths: A Formalized Approach to Computing Fundamental Groups

Abstract.
The Seifert-van Kampen theorem computes the fundamental group of a space from the fundamental groups of its constituents. We develop a modular SVK framework within the setting of computational paths—an approach to equality where witnesses are explicit sequences of rewrites governed by the $\mathrm{LND}_{\mathrm{EQ}}$-TRS.
Our contributions are: (i) pushouts as higher-inductive types with modular typeclass assumptions for computation rules; (ii) free products and amalgamated free products as quotients of word representations; (iii) an SVK equivalence schema $\pi_{1}(\mathrm{Pushout}(A,B,C))\simeq\pi_{1}(A)\ast_{\pi_{1}(C)}\pi_{1}(B)$ parametric in user-supplied encode/decode structure—the decode map and amalgamation-respecting properties are proved, while the encode map and round-trip properties are typeclass assumptions; and (iv) instantiations for classical spaces: figure-eight ($\pi_{1}(S^{1}\vee S^{1})\simeq\mathbb{Z}\ast\mathbb{Z}$), spheres ($\pi_{1}(S^{n})\simeq 1$ for $n\geq 2$), Klein bottle, real projective plane, orientable/non-orientable surfaces, and lens spaces.
The development is formalized in Lean 4 with 41,130 lines across 107 modules, using 36 kernel axioms for HIT type/constructor declarations. Pushout computation rules are not kernel axioms—they are typeclass assumptions, making dependencies explicit. The library also includes extended features (higher homotopy groups, covering spaces, fibrations, Mayer-Vietoris) documented in Section 9.
Key words and phrases:
Seifert-van Kampen theorem, computational paths, fundamental group, higher-inductive types, pushouts, free products, type theory
2020 Mathematics Subject Classification:
Primary: 55Q05; Secondary: 03B38, 18N10, 68V15
1. Introduction
The Seifert-van Kampen theorem, first proved by Herbert Seifert [16] and Egbert van Kampen [18], is one of the most powerful tools in algebraic topology for computing fundamental groups (see [7, 11] for modern treatments). It states that if a path-connected space $X$ is the union of two path-connected open subspaces $U$ and $V$ with path-connected intersection $U\cap V$, then the fundamental group $\pi_{1}(X)$ can be computed as the amalgamated free product:
$\pi_{1}(X)\cong\pi_{1}(U)\ast_{\pi_{1}(U\cap V)}\pi_{1}(V)$
In homotopy type theory (HoTT) [17], this theorem takes a particularly elegant form when expressed in terms of pushouts—a higher-inductive type (HIT) that generalizes the notion of gluing spaces together. The HoTT version of the Seifert-van Kampen theorem was developed by Favonia and Shulman [8], providing a synthetic proof that works entirely within type theory.
The present paper develops a modular SVK framework within the setting of computational paths [2, 12]—an alternative approach to equality in type theory where witnesses of equality are explicit sequences of rewrites. This approach offers several advantages:
  1. (1)
    Explicit witnesses: Every equality proof carries the full computational content showing why two terms are equal.
  2. (2)
    Syntactic path equality: The rewrite equality relation ($\sim$) is defined as the symmetric-transitive closure of Step rules. Under termination and confluence (established in prior work [14], not fully mechanized here), $\sim$ is decidable via normalization.
  3. (3)
    Concrete coherence: Higher coherence witnesses (pentagon, triangle, etc.) are explicit rewrite derivations rather than abstract existence proofs.
  4. (4)
    Modular assumptions: The SVK framework uses split typeclass assumptions, allowing users to provide only what they need—from full encode/decode data to minimal Prop-level bijectivity claims.
Relation to Classical SVK.
The classical theorem requires $X=U\cup V$ with $U,V,U\cap V$ open and path-connected. The HIT version generalizes this: the pushout $\mathrm{Pushout}(A,B,C,f,g)$ corresponds to the union where $f$ and $g$ specify how the “intersection” $C$ embeds into both spaces. Path-connectedness assumptions are needed to ensure the fundamental group is well-defined; in the HIT setting, this becomes the assumption that all types are path-connected.
1.1. Main Contributions
This paper provides:
  1. (1)
    Pushouts implemented via Lean’s quotient types (zero kernel axioms), with modular typeclass assumptions for computation rules:
    • Point constructors $\mathrm{inl}:A\to\mathrm{Pushout}$ and $\mathrm{inr}:B\to\mathrm{Pushout}$ (definitions, not axioms)
    • Path constructor $\mathrm{glue}:\prod_{c:C}\mathrm{Path}(\mathrm{inl}(f(c)),\mathrm{inr}(g(c)))$ (via Quot.sound)
    • Computation/naturality hypotheses as typeclasses (HasRecGlueRwEq, HasGlueNaturalLoopRwEq)
    • Automatic instance derivation for common cases (Axiom K, decidable equality, subsingletons)
  2. (2)
    The construction of free products $G_{1}\ast G_{2}$ and amalgamated free products $G_{1}\ast_{H}G_{2}$, including:
    • Standard amalgamation quotient
    • Full amalgamated free product with free group reduction (FullAmalgamatedFreeProduct)
    • Both data-level and Prop-level (choice-based) interfaces
  3. (3)
    An SVK equivalence schema parametric in user-supplied encode/decode structure, with split typeclass assumptions:
    • HasPushoutSVKEncodeQuot: encode map (assumed)
    • HasPushoutSVKDecodeEncode: decode-encode round-trip (assumed)
    • HasPushoutSVKEncodeDecode: encode-decode round-trip (assumed)
    • HasPushoutSVKDecodeAmalgBijective: Prop-level bijectivity alternative
    • Decode map and amalgamation-respecting property: proved
  4. (4)
    Applications demonstrating the power of the SVK framework:
    • The wedge sum: $\pi_{1}(A\vee B)\simeq\pi_{1}(A)\ast\pi_{1}(B)$
    • The figure-eight: $\pi_{1}(S^{1}\vee S^{1})\simeq\mathbb{Z}\ast\mathbb{Z}$
    • Spheres: $\pi_{1}(S^{n})\simeq 1$ for $n\geq 2$
    • Klein bottle: $\pi_{1}(K)\simeq\mathbb{Z}\rtimes\mathbb{Z}$ (semidirect product)
    • Real projective plane: $\pi_{1}(\mathbb{RP}^{2})\simeq\mathbb{Z}_{2}$
    • Orientable surfaces: $\pi_{1}(\Sigma_{g})\simeq\langle a_{1},b_{1},\ldots,a_{g},b_{g}\mid\prod_{i=1}^{g}[a_{i},b_{i}]=1\rangle$
    • Non-orientable surfaces: $\pi_{1}(N_{n})\simeq\langle a_{1},\ldots,a_{n}\mid a_{1}^{2}\cdots a_{n}^{2}=1\rangle$
    • Lens spaces: $\pi_{1}(L(p,q))\simeq\mathbb{Z}/p\mathbb{Z}$ for coprime $p,q$
    • Bouquet of $n$ circles: $\pi_{1}(\bigvee_{n}S^{1})\simeq F_{n}$ (free group)
  5. (5)
    A complete Lean 4 formalization comprising:
    • 41,130 lines of code across 107 modules
    • Only 36 kernel axioms (all for HITs)
    • Torus constructed as $S^{1}\times S^{1}$ (no axioms)
    • Opt-in axiom imports for assumption-free APIs
    • Automated path simplification tactics
1.2. Related Work
The Seifert-van Kampen theorem in HoTT was first formalized by Favonia and Shulman [8], who developed a general framework for computing fundamental groups of HITs. Their work uses the encode-decode method pioneered by Licata and Shulman [10] for computing $\pi_{1}(S^{1})\cong\mathbb{Z}$.
The computational paths approach to equality originates in work by de Queiroz and colleagues [2, 12, 13], building on earlier ideas about equality proofs as sequences of rewrites [3]. Our previous work established that computational paths form a weak groupoid [6] and can be used to calculate fundamental groups [4, 5].
The connection between rewriting and coherence has been developed by Kraus and von Raumer [9], who showed that confluence and termination suffice to build coherent higher-dimensional structure. Their key insight is that in a confluent, terminating rewrite system, any two parallel reductions to normal form can be connected by a canonical “diamond” of rewrite paths—and this extends to higher dimensions. For our $\mathrm{LND}_{\mathrm{EQ}}$-TRS, this means that higher coherences (such as the pentagon identity for associativity and the naturality conditions for transport) are automatically satisfied once we establish that paths with the same normal form are $\sim$-equivalent.
1.3. Reading Guide
Readers interested in the mathematical content can focus on Sections 36, treating Lean code as pseudocode. Those interested in formalization should additionally consult Section 8 and the supplementary code. Key definitions are boxed for quick reference. We use the figure-eight space $S^{1}\vee S^{1}$ as a running example throughout.
1.4. Structure of the Paper
Section 2 reviews computational paths, rewrite equality, and fundamental groups. Section 3 defines pushouts as higher-inductive types with modular assumptions. Section 4 constructs free products and amalgamated free products. Section 5 proves the Seifert-van Kampen theorem using the encode-decode method with split assumptions. Section 6 applies the theorem extensively. Section 7 covers extended features including higher homotopy groups and covering spaces. Section 8 discusses the Lean 4 formalization. Section 10 concludes.
2. Background: Computational Paths
2.1. Computational Paths
A computational path $p:\mathrm{Path}(a,b)$ from $a$ to $b$ (both terms of type $A$) is an explicit sequence of rewrite steps witnessing the equality $a=b$. The fundamental operations are:
  • Reflexivity: $\rho(a):\mathrm{Path}(a,a)$ — the empty sequence of rewrites.
  • Symmetry: If $p:\mathrm{Path}(a,b)$, then $\sigma(p):\mathrm{Path}(b,a)$ — reverse and invert each step.
  • Transitivity: If $p:\mathrm{Path}(a,b)$ and $q:\mathrm{Path}(b,c)$, then $p\cdot q:\mathrm{Path}(a,c)$ — concatenate the sequences.
  • Congruence: If $p:\mathrm{Path}(a,b)$ and $f:A\to B$, then $f_{*}(p):\mathrm{Path}(f(a),f(b))$.
  • Transport: If $p:\mathrm{Path}(a,b)$ and $P:A\to\mathrm{Type}$, then $\mathrm{transport}(P,p):P(a)\to P(b)$.
Notation.
Throughout this paper, we use:
  • $p\cdot q$ for path composition (transitivity)
  • $p^{-1}$ for path inverse (symmetry)
  • $f_{*}(p)$ for $\mathrm{congrArg}(f,p)$
  • $\rho(a)$ or simply $\rho$ for reflexivity
In Lean, a path is represented as a structure storing both an explicit rewrite trace and an equality proof:
1structure Path {A : Type u} (a b : A) where
2 steps : List (Step A) -- explicit list of rewrite steps
3 proof : a = b -- the derived equality proof
Remark 2.1 (Explicit Step Lists and Consistency).
The steps field is the distinguishing data of a path, not the underlying equality proof. Two paths with the same endpoints may have different step lists, and they are related by $\sim$ only if there is an explicit derivation transforming one step sequence into the other via the Step rules.
This design is critical for consistency: in Lean’s proof-irrelevant Prop, all equality proofs $p,q:a=b$ satisfy $p=q$ (proof irrelevance). If paths were identified solely by their equality proofs, all loops at a point would be equal, collapsing $\pi_{1}$ to the trivial group.
Instead, path identity depends on the explicit steps list. The fundamental group $\pi_{1}(A,a):=\mathrm{Loop}(A,a)/{\sim}$ quotients loops by rewrite equivalence of their step sequences—not by equality of their underlying proofs. The $\sim$ relation captures the computational content: two loops are identified only when one step sequence can be transformed to the other via the groupoid laws.
2.2. The Step Relation
The foundation of the computational paths framework is the Step relation, which defines single-step rewrites between paths. The $\mathrm{LND}_{\mathrm{EQ}}$-TRS consists of over 50 rewrite rules organized into categories:
  • Groupoid laws: $\rho^{-1}\rightsquigarrow\rho$, $(p^{-1})^{-1}\rightsquigarrow p$, $\rho\cdot p\rightsquigarrow p$, $\cdot\rho\rightsquigarrow p, pp$$\cdot p^{-1}\rightsquigarrow\rho$, $p^{-1}\cdot p\rightsquigarrow\rho$, $(p\cdot q)^{-1}\rightsquigarrow q^{-1}\cdot p^{-1}$, and $(p\cdot q)\cdot r\rightsquigarrow p\cdot(q\cdot r)$.
  • Type-specific rules: $\beta$-rules for products, sums, and functions; $\eta$-rules; transport laws.
  • Context rules: Allow rewrites inside larger expressions: if $p\rightsquigarrow q$ then $C[p]\rightsquigarrow C[q]$.
  • Congruence closure: If $p\rightsquigarrow q$ then $p^{-1}\rightsquigarrow q^{-1}$, $p\cdot r\rightsquigarrow q\cdot r$, and $r\cdot p\rightsquigarrow r\cdot q$.
In Lean, the Step relation is defined inductively with 76 constructors:
1inductive Step : {A : Type u} -> {a b : A} -> Path a b -> Path a b -> Prop
2 | symm_refl (a : A) : Step (symm (Path.refl a)) (Path.refl a)
3 | symm_symm (p : Path a b) : Step (symm (symm p)) p
4 | trans_refl_left (p : Path a b) : Step (trans (Path.refl a) p) p
5 | trans_refl_right (p : Path a b) : Step (trans p (Path.refl b)) p
6 | trans_symm (p : Path a b) : Step (trans p (symm p)) (Path.refl a)
7 | symm_trans (p : Path a b) : Step (trans (symm p) p) (Path.refl b)
8 | trans_assoc (p : Path a b) (q : Path b c) (r : Path c d) :
9 Step (trans (trans p q) r) (trans p (trans q r))
10 | symm_congr : Step p q -> Step (symm p) (symm q)
11 | trans_congr_left (r : Path b c) : Step p q -> Step (trans p r) (trans q r)
12 | trans_congr_right (p : Path a b) : Step q r -> Step (trans p q) (trans p r)
13 -- ... plus type-specific rules (products, sums, transport, contexts)
2.3. Rewrite Equality
Two paths $p,q:\mathrm{Path}(a,b)$ are rewrite equal ($p\sim q$) if they can be transformed into each other via the $\mathrm{LND}_{\mathrm{EQ}}$-TRS rewrite system. The key rewrite rules include:
$\rho\cdot p$ $\rightsquigarrow p$ (left unit)
$p\cdot\rho$ $\rightsquigarrow p$ (right unit)
$p\cdot p^{-1}$ $\rightsquigarrow\rho$ (right inverse)
$p^{-1}\cdot p$ $\rightsquigarrow\rho$ (left inverse)
$(p\cdot q)\cdot r$ $\rightsquigarrow p\cdot(q\cdot r)$ (associativity)
Example 2.2 (Rewrite Derivation).
Consider $p\cdot(q\cdot q^{-1})$ where $p,q:\mathrm{Path}(a,a)$. The derivation to $p$ proceeds:
$p\cdot(q\cdot q^{-1})$ $\rightsquigarrow p\cdot\rho$ (right inverse)
$\rightsquigarrow p$ (right unit)
This explicit derivation is the “computational content” witnessing that $p\cdot(q\cdot q^{-1})\sim p$.
In Lean, rewrite equality is defined as the equivalence closure of Step:
1inductive RwEq {A : Type u} {a b : A} : Path a b -> Path a b -> Prop
2 | refl (p : Path a b) : RwEq p p
3 | step {p q : Path a b} : Step p q -> RwEq p q
4 | symm {p q : Path a b} : RwEq p q -> RwEq q p
5 | trans {p q r : Path a b} : RwEq p q -> RwEq q r -> RwEq p r
2.4. Metatheory of the $\mathrm{LND}_{\mathrm{EQ}}$-TRS
The rewrite system $\mathrm{LND}_{\mathrm{EQ}}$-TRS has been studied in prior work [13, 14, 2]. The following table summarizes the status of key metatheoretic properties in this formalization:
Property Status Location / Notes
Soundness Proved step_toEq in Step.lean: $p\sim q\Rightarrow p.\mathrm{proof}=q.\mathrm{proof}$
Confluence Typeclass HasJoinOfRw in Confluence.lean: users provide join witnesses
Local confluence Proved Critical pairs analyzed in Confluence.lean (specific cases)
Termination Not formalized Prior paper proofs [14]; not mechanized here
Decidability of $\sim$ Assumed Would follow from termination + confluence; not fully mechanized
Remark 2.3 (Non-symmetric Soundness).
Soundness says $p\sim q\Rightarrow p.\mathrm{proof}=q.\mathrm{proof}$. Critically, the converse does not hold: paths with the same equality proof are not necessarily $\sim$-related. This asymmetry is essential for non-trivial fundamental groups (see Remark 2.1).
2.5. Loop Spaces and Fundamental Groups
Definition 2.4 (Path-Connected).
A type $A$ is path-connected if for all $a,b:A$, there exists $p:\mathrm{Path}(a,b)$. Equivalently, $\prod_{a,b:A}\|\mathrm{Path}(a,b)\|_{-1}$ where $\|-\|_{-1}$ denotes propositional truncation.
Definition 2.5 (Loop Space).
The loop space of a type $A$ at a basepoint $a:A$ is:
$\Omega(A,a):=\mathrm{Path}(a,a)$
Definition 2.6 (Fundamental Group).
The fundamental group $\pi_{1}(A,a)$ is the quotient of the loop space by rewrite equality:
$\pi_{1}(A,a):=\Omega(A,a)/{\sim}$
In Lean, the fundamental group is realized as a quotient:
1abbrev LoopSpace (A : Type u) (a : A) : Type u := Path a a
2
3def PiOne (A : Type u) (a : A) : Type u := Quot (@RwEq A a a)
4
5notation "pi_1(" A "," a ")" => PiOne A a
The group operations on $\pi_{1}(A,a)$ are induced from path operations:
  • Identity: $\rho]$
  • Multiplication: $\alpha]\cdot[\beta]:=[\alpha\cdot\beta]$
  • Inverse: $\alpha]^{-1}:=[\alpha^{-1}]$
The group laws (associativity, unit, inverse) hold because the corresponding path identities are witnessed by $\sim$.
Remark 2.7 (Base Point Dependence).
The fundamental group $\pi_{1}(A,a)$ depends on the choice of basepoint $a$. For different basepoints $a,b:A$ in a path-connected type, the fundamental groups are isomorphic via conjugation by a path $\gamma:\mathrm{Path}(a,b)$:
$\pi_{1}(A,a)\cong\pi_{1}(A,b),\quad[\alpha]\mapsto[\gamma^{-1}\cdot\alpha\cdot\gamma]$
However, this isomorphism is not canonical—it depends on the choice of $\gamma$.
2.6. The Circle and Its Fundamental Group
Definition 2.8 (Circle HIT).
The circle $S^{1}$ is the higher-inductive type with:
  • A point constructor: $\mathrm{base}:S^{1}$
  • A path constructor: $\mathrm{loop}:\mathrm{Path}(\mathrm{base},\mathrm{base})$
Theorem 2.9 (Fundamental Group of Circle).
$\pi_{1}(S^{1},\mathrm{base})\simeq\mathbb{Z}$
This classical result, first formalized via encode-decode by Licata and Shulman [10], serves as the foundation for our SVK applications.
3. Pushouts as Higher-Inductive Types
3.1. The Pushout Type
Given types $A$, $B$, $C$ with maps $f:C\to A$ and $g:C\to B$, the pushout $\mathrm{Pushout}(A,B,C,f,g)$ is the higher-inductive type that “glues” $A$ and $B$ together along the common image of $C$:
[…]
Definition 3.1 (Pushout).
The pushout $\mathrm{Pushout}(A,B,C,f,g)$ is characterized by:
  1. (1)
    Point constructors:
    $\mathrm{inl}$ $:A\to\mathrm{Pushout}(A,B,C,f,g)$
    $\mathrm{inr}$ $:B\to\mathrm{Pushout}(A,B,C,f,g)$
  2. (2)
    Path constructor: For each $c:C$,
    $\mathrm{glue}(c):\mathrm{Path}(\mathrm{inl}(f(c)),\mathrm{inr}(g(c)))$
In Lean, pushouts are implemented using quotient types (no kernel axioms):
1-- Relation generating the pushout quotient
2inductive PushoutRel (A B C : Type u) (f : C -> A) (g : C -> B)
3 : Sum A B -> Sum A B -> Prop
4 | glue (c : C) : PushoutRel A B C f g (Sum.inl (f c)) (Sum.inr (g c))
5
6-- Pushout as quotient of Sum by PushoutRel
7def Pushout (A B C : Type u) (f : C -> A) (g : C -> B) : Type u :=
8 Quot (PushoutRel A B C f g)
9
10-- Constructors are definitions, not axioms
11def inl (a : A) : Pushout A B C f g := Quot.mk _ (Sum.inl a)
12def inr (b : B) : Pushout A B C f g := Quot.mk _ (Sum.inr b)
13def glue (c : C) : Path (inl (f c)) (inr (g c)) :=
14 Path.ofEq (Quot.sound (PushoutRel.glue c))
Remark 3.2 (Quot-Based Implementation).
Unlike Circle, Klein bottle, and other HITs which require kernel axioms for their path constructors, the Pushout uses Lean’s built-in Quot type. This adds zero kernel axioms to the trusted base. The glue path is constructed via Quot.sound, which witnesses that related elements become equal in the quotient.
3.2. Recursion Principle
Definition 3.3 (Pushout Recursion).
Given a type $D$ with:
  • $\mathrm{onInl}:A\to D$
  • $\mathrm{onInr}:B\to D$
  • $\mathrm{onGlue}:\prod_{c:C}\mathrm{Path}(\mathrm{onInl}(f(c)),\mathrm{onInr}(g(c)))$
There exists a function $\mathrm{rec}:\mathrm{Pushout}(A,B,C,f,g)\to D$ such that:
$\mathrm{rec}(\mathrm{inl}(a))$ $=\mathrm{onInl}(a)$
$\mathrm{rec}(\mathrm{inr}(b))$ $=\mathrm{onInr}(b)$
$(\mathrm{rec})_{*}(\mathrm{glue}(c))$ $\sim\mathrm{onGlue}(c)$
3.3. Modular Typeclass Assumptions
A key design choice in our formalization is to make the computation and naturality rules typeclass assumptions rather than kernel axioms. This allows users to provide only what they need:
1class HasRecGlueRwEq (D : Type u) (onInl : A -> D) (onInr : B -> D)
2 (onGlue : forall c, Path (onInl (f c)) (onInr (g c))) where
3 rec_glue_rweq : forall c, RwEq (congrArg rec (glue c)) (onGlue c)
4
5class HasGlueNaturalLoopRwEq (c0 : C) where
6 glue_natural_loop : forall (p : LoopSpace C c0),
7 RwEq (inlPath (congrArg f p))
8 (trans (glue c0) (trans (inrPath (congrArg g p)) (symm (glue c0))))
3.4. Automatic Instance Derivation
The framework provides automatic instances for common cases:
1-- When both A and B satisfy Axiom K
2instance [DecidableEq A] [HasDecidableEqAxiomK A]
3 [DecidableEq B] [HasDecidableEqAxiomK B] :
4 HasGlueNaturalLoopRwEq c0 where
5 glue_natural_loop := fun p => by path_simp
6
7-- When C is a subsingleton (e.g., Unit for wedge sums)
8instance [Subsingleton C] : HasGlueNaturalLoopRwEq c0 where
9 glue_natural_loop := fun p => by simp [Subsingleton.elim]; path_rfl
3.5. Glue Naturality
A key property for the SVK theorem is the naturality of glue paths:
Lemma 3.4 (Glue Naturality).
For any path $p:\mathrm{Path}(c_{1},c_{2})$ in $C$, the following square commutes up to $\sim$:
$\mathrm{inl}(f(c_{1}))$$\mathrm{inl}(f(c_{2}))$$\mathrm{inr}(g(c_{1}))$$\mathrm{inr}(g(c_{2}))$$\scriptstyle{\mathrm{inl}_{*}(f_{*}(p))}$$\scriptstyle{\mathrm{glue}(c_{1})}$$\scriptstyle{\mathrm{glue}(c_{2})}$$\scriptstyle{\mathrm{inr}_{*}(g_{*}(p))}$
That is, we have the rewrite equality:
$\mathrm{inl}_{*}(f_{*}(p))\cdot\mathrm{glue}(c_{2})\sim\mathrm{glue}(c_{1})\cdot\mathrm{inr}_{*}(g_{*}(p))$
Rearranging, we obtain a form useful for the decode function:
Corollary 3.5.
For any path $p:\mathrm{Path}(c_{1},c_{2})$ in $C$:
$\mathrm{inl}_{*}(f_{*}(p))\sim\mathrm{glue}(c_{1})\cdot\mathrm{inr}_{*}(g_{*}(p))\cdot\mathrm{glue}(c_{2})^{-1}$
Remark 3.6 (Two Forms of Glue Naturality).
The axiom glue_natural provides a propositional equality of paths, while glue_natural_rweq provides the $\sim$ equivalence needed for quotient arguments. In a cubical type theory, both would be provable; here, we axiomatize them.
3.6. Special Cases
3.6.1. Wedge Sum
The wedge sum $A\vee B$ is the pushout of $A\leftarrow\mathrm{pt}\to B$:
$A\vee B:=\mathrm{Pushout}(A,B,\mathbf{1},\lambda\_.a_{0},\lambda\_.b_{0})$
This identifies the basepoints $a_{0}:A$ and $b_{0}:B$.
Example 3.7 (Figure-Eight as Wedge).
The figure-eight space is $S^{1}\vee S^{1}$—two circles joined at a single point. The glue path identifies the basepoints of the two circles:
$\mathrm{glue}(*):\mathrm{Path}(\mathrm{inl}(\mathrm{base}_{1}),\mathrm{inr}(\mathrm{base}_{2}))$
In Lean:
1def Wedge (A : Type u) (B : Type u) (a0 : A) (b0 : B) : Type u :=
2 Pushout A B PUnit (fun _ => a0) (fun _ => b0)
3.6.2. Suspension
The suspension $\Sigma A$ is the pushout of $\mathbf{1}\leftarrow A\to\mathbf{1}$:
$\Sigma A:=\mathrm{Pushout}(\mathbf{1},\mathbf{1},A,\lambda\_.*,\lambda\_.*)$
This adds a “north pole”, “south pole”, and meridian paths connecting them.
In Lean:
1def Suspension (A : Type u) : Type u :=
2 Pushout PUnit PUnit A (fun _ => PUnit’.unit) (fun _ => PUnit’.unit)
3
4noncomputable def north : Suspension A := Pushout.inl PUnit’.unit
5noncomputable def south : Suspension A := Pushout.inr PUnit’.unit
6noncomputable def merid (a : A) : Path north south := Pushout.glue a
3.6.3. Bouquet of $n$ Circles
The bouquet of $n$ circles $\bigvee_{n}S^{1}$ generalizes the wedge sum:
1def BouquetN (n : Nat) : Type u :=
2 -- n circles wedged together at a single point
3 Pushout (Fin n -> Circle) PUnit (Fin n)
4 (fun i => Circle.circleBase)
5 (fun _ => PUnit’.unit)
4. Free Products and Amalgamated Free Products
4.1. Free Product Words
Definition 4.1 (Free Product Word).
A word in the free product $G_{1}\ast G_{2}$ is an alternating sequence of elements:
1inductive FreeProductWord (G1 : Type u) (G2 : Type u) : Type u
2 | nil : FreeProductWord G1 G2
3 | consLeft (x : G1) (rest : FreeProductWord G1 G2)
4 | consRight (y : G2) (rest : FreeProductWord G1 G2)
The group operations on words are defined by concatenation and inversion with appropriate reductions.
Definition 4.2 (Word Operations).
1-- Concatenation
2def concat : FreeProductWord G1 G2 -> FreeProductWord G1 G2 ->
3 FreeProductWord G1 G2
4 | nil, w => w
5 | consLeft x rest, w => consLeft x (concat rest w)
6 | consRight y rest, w => consRight y (concat rest w)
7
8-- Inversion
9def invert : FreeProductWord G1 G2 -> FreeProductWord G1 G2
10 | nil => nil
11 | consLeft x rest => concat (invert rest) (singleLeft (-x))
12 | consRight y rest => concat (invert rest) (singleRight (-y))
4.2. Amalgamated Free Product
When $G_{1}$ and $G_{2}$ share a common subgroup $H$ via homomorphisms $i_{1}:H\to G_{1}$ and $i_{2}:H\to G_{2}$, the amalgamated free product $G_{1}\ast_{H}G_{2}$ identifies $i_{1}(h)$ with $i_{2}(h)$ for all $h:H$.
Definition 4.3 (Amalgamation Relation).
1inductive AmalgRelation (i1 : H -> G1) (i2 : H -> G2) :
2 FreeProductWord G1 G2 -> FreeProductWord G1 G2 -> Prop
3 | amalgLeftToRight (h : H) (pre suf : FreeProductWord G1 G2) :
4 AmalgRelation i1 i2
5 (concat pre (concat (singleLeft (i1 h)) suf))
6 (concat pre (concat (singleRight (i2 h)) suf))
Definition 4.4 (Amalgamated Free Product).
1def AmalgEquiv (i1 : H -> G1) (i2 : H -> G2) :=
2 EqvGen (AmalgRelation i1 i2)
3
4def AmalgamatedFreeProduct (G1 G2 H : Type u)
5 (i1 : H -> G1) (i2 : H -> G2) : Type u :=
6 Quot (AmalgEquiv i1 i2)
4.3. Full Amalgamated Free Product
For a “canonical” target with fully reduced words, we define the full amalgamated free product that additionally reduces via free group laws:
1inductive FreeGroupStep : FreeProductWord G1 G2 -> FreeProductWord G1 G2 -> Prop
2 | cancel_left (x : G1) (pre suf : FreeProductWord G1 G2)
3 (h : x + (-x) = 0) :
4 FreeGroupStep (concat pre (consLeft x (consLeft (-x) suf)))
5 (concat pre suf)
6 | cancel_right (y : G2) (pre suf : FreeProductWord G1 G2)
7 (h : y + (-y) = 0) :
8 FreeGroupStep (concat pre (consRight y (consRight (-y) suf)))
9 (concat pre suf)
10
11def FullAmalgEquiv := EqvGen (fun w1 w2 =>
12 AmalgRelation i1 i2 w1 w2 \/ FreeGroupStep w1 w2)
13
14def FullAmalgamatedFreeProduct (G1 G2 H : Type u) (i1 : H -> G1) (i2 : H -> G2) :=
15 Quot (FullAmalgEquiv i1 i2)
Remark 4.5 (Standard vs Full Target).
The AmalgamatedFreeProduct is the standard target where words are only quotiented by amalgamation. The FullAmalgamatedFreeProduct additionally applies free group reductions. Both are valid SVK targets; the choice depends on application needs.
5. The Seifert-van Kampen Theorem
5.1. Statement
Theorem 5.1 (Seifert-van Kampen Schema).
Let $A$, $B$, $C$ be path-connected types with maps $f:C\to A$ and $g:C\to B$, and let $c_{0}:C$. Then:
$\pi_{1}(\mathrm{Pushout}(A,B,C,f,g),\mathrm{inl}(f(c_{0})))\simeq\pi_{1}(A,f(c_{0}))\ast_{\pi_{1}(C,c_{0})}\pi_{1}(B,g(c_{0}))$
Dependency Summary for Theorem 5.1 Proved: The decode map $\mathrm{decode}:\mathrm{Word}\to\pi_{1}(\mathrm{Pushout})$ is constructed explicitly. Proved: Decode respects the amalgamation relation (Lemma 5.2), assuming glue naturality (HasGlueNaturalLoopRwEq). Proved: The quotient map $\mathrm{decode}_{\mathrm{Amalg}}:\mathrm{AmalgFP}\to\pi_{1}(\mathrm{Pushout})$ is well-defined. $\square$ Assumed (typeclass): The encode map exists (HasPushoutSVKEncodeQuot). $\square$ Assumed (typeclass): Decode-encode round-trip (HasPushoutSVKDecodeEncode). $\square$ Assumed (typeclass): Encode-decode round-trip (HasPushoutSVKEncodeDecode). The equivalence is therefore parametric in user-supplied encode/decode structure. Alternatively, the choice-based interface (HasPushoutSVKDecodeAmalgBijective) requires only a Prop-level bijectivity assumption without explicit encode.
5.2. Split Assumptions Architecture
The SVK proof is structured around split typeclass assumptions, allowing maximum flexibility:
1-- Just the encode map
2class HasPushoutSVKEncodeQuot (c0 : C) where
3 pushoutEncodeQuot : pi_1(Pushout A B C f g, inl (f c0)) ->
4 AmalgamatedFreeProduct (pi_1 A) (pi_1 B) (pi_1 C)
5
6-- Decode-encode round-trip (Prop)
7class HasPushoutSVKDecodeEncode (c0 : C) where
8 decode_encode : forall alpha, pushoutDecodeAmalg c0 (encode alpha) = alpha
9
10-- Encode-decode round-trip (up to AmalgEquiv)
11class HasPushoutSVKEncodeDecode (c0 : C) where
12 encode_decode : forall w, AmalgEquiv (encode (decode w)) w
13
14-- Encode-decode up to FullAmalgEquiv
15class HasPushoutSVKEncodeDecodeFull (c0 : C) where
16 encode_decode_full : forall w, FullAmalgEquiv (encode (decode w)) w
17
18-- Prop-level bijectivity (choice-based)
19class HasPushoutSVKDecodeAmalgBijective (c0 : C) where
20 decode_injective : Function.Injective (pushoutDecodeAmalg c0)
21 decode_surjective : Function.Surjective (pushoutDecodeAmalg c0)
5.3. Multiple Interfaces
The framework provides multiple equivalences depending on available assumptions:
1-- Standard SVK with explicit encode data
2noncomputable def seifertVanKampenEquiv
3 [HasPushoutSVKEncodeQuot c0] [HasGlueNaturalLoopRwEq c0]
4 [HasPushoutSVKDecodeEncode c0] [HasPushoutSVKEncodeDecode c0] :
5 SimpleEquiv (pi_1(Pushout A B C f g, inl (f c0)))
6 (AmalgamatedFreeProduct (pi_1 A) (pi_1 B) (pi_1 C))
7
8-- SVK with full target (free reduction)
9noncomputable def seifertVanKampenFullEquiv
10 [HasPushoutSVKEncodeQuot c0] [HasGlueNaturalLoopRwEq c0]
11 [HasPushoutSVKDecodeEncode c0] [HasPushoutSVKEncodeDecodeFull c0] :
12 SimpleEquiv (pi_1(Pushout A B C f g, inl (f c0)))
13 (FullAmalgamatedFreeProduct (pi_1 A) (pi_1 B) (pi_1 C))
14
15-- Choice-based SVK (no explicit encode required)
16noncomputable def seifertVanKampenEquiv_of_decodeAmalg_bijective
17 [HasPushoutSVKDecodeAmalgBijective c0] [HasGlueNaturalLoopRwEq c0] :
18 SimpleEquiv (pi_1(Pushout A B C f g, inl (f c0)))
19 (AmalgamatedFreeProduct (pi_1 A) (pi_1 B) (pi_1 C))
5.4. The Encode-Decode Method
5.4.1. The Decode Map
The decode map converts a word to a loop in the pushout:
1noncomputable def pushoutDecode (c0 : C) :
2 FreeProductWord (pi_1(A, f c0)) (pi_1(B, g c0))
3 -> pi_1(Pushout A B C f g, inl (f c0))
4 | .nil => Quot.mk _ (Path.refl _)
5 | .consLeft alpha rest =>
6 piOneMul (liftLeft alpha) (pushoutDecode c0 rest)
7 | .consRight beta rest =>
8 piOneMul (conjugateRight beta) (pushoutDecode c0 rest)
where conjugateRight conjugates by the glue path:
$\beta\mapsto\mathrm{glue}(c_{0})\cdot\mathrm{inr}_{*}(\beta)\cdot\mathrm{glue}(c_{0})^{-1}$
5.4.2. Decode Respects Amalgamation
Lemma 5.2 (Decode Respects Amalgamation).
For any $\gamma:\pi_{1}(C,c_{0})$:
$\mathrm{decode}(\mathrm{consLeft}(f_{*}(\gamma),w))=\mathrm{decode}(\mathrm{consRight}(g_{*}(\gamma),w))$
Proof.
This follows from glue naturality: $\mathrm{inl}_{*}(f_{*}(p))\sim\mathrm{glue}(c_{0})\cdot\mathrm{inr}_{*}(g_{*}(p))\cdot\mathrm{glue}(c_{0})^{-1}$. ∎
In Lean:
1theorem pushoutDecodeAmalg_respects_rel [HasGlueNaturalLoopRwEq c0]
2 (h : HasAmalgRelation i1 i2 w1 w2) :
3 pushoutDecodeAmalg c0 w1 = pushoutDecodeAmalg c0 w2 := by
4 induction h with
5 | amalgLeftToRight gamma pre suf =>
6 -- Use glue naturality to show the images are equal
7 simp only [pushoutDecodeAmalg]
8 apply congrArg
9 -- Key step: glue_natural_loop shows the conjugation
10 exact glue_natural_loop_pi1 gamma
5.4.3. The Encode Map
The encode map is axiomatized via the typeclass HasPushoutSVKEncodeQuot. In a cubical type theory, this would be constructed via the code family method.
5.4.4. Round-Trip Properties
The round-trip properties establish the equivalence:
  • $\mathrm{decode}(\mathrm{encode}(\alpha))=\alpha$ (via HasPushoutSVKDecodeEncode)
  • $\mathrm{encode}(\mathrm{decode}(w))\sim w$ up to amalgamation (via HasPushoutSVKEncodeDecode)
6. Applications
We present $\pi_{1}$ calculations using several methods. Table 1 summarizes the results, distinguishing between direct SVK applications (pushouts with trivial or simple amalgamation), CW-complex constructions (2-cells attached via SVK), and other methods (products, fibrations).
Table 1. Summary of $\pi_{1}$ calculations
Space Fundamental Group Method SVK?
Direct SVK (pushouts with simple amalgamation)
$A\vee B$ $\pi_{1}(A)\ast\pi_{1}(B)$ Pushout over $\mathbf{1}$ Yes
$S^{1}\vee S^{1}$ $\mathbb{Z}\ast\mathbb{Z}$ Wedge of circles Yes
$\bigvee_{n}S^{1}$ $F_{n}$ (free group) $n$-bouquet pushout Yes
$S^{n}$ ($n\geq 2$) 11 Suspension pushout Yes
CW-complex via SVK (2-cell attachment)
$K$ (Klein bottle) $\mathbb{Z}\rtimes\mathbb{Z}$ $S^{1}\vee S^{1}\cup_{aba^{-1}b}D^{2}$ Yes
$\mathbb{RP}^{2}$ $\mathbb{Z}_{2}$ $S^{1}\cup_{a^{2}}D^{2}$ Yes
$\Sigma_{g}$ (genus $g$) Surface group $\bigvee_{2g}S^{1}\cup_{\prod[a_{i},b_{i}]}D^{2}$ Yes
$N_{n}$ (crosscap $n$) $\langle a_{i}\mid\prod a_{i}^{2}=1\rangle$ $\bigvee_{n}S^{1}\cup_{\prod a_{i}^{2}}D^{2}$ Yes
$L(p,q)$ $\mathbb{Z}/p\mathbb{Z}$ Heegaard (tori gluing) Yes
Other methods (not SVK)
$T^{2}$ (torus) $\mathbb{Z}\times\mathbb{Z}$ Product: $\pi_{1}(A\times B)$ No
$\mathbb{CP}^{n}$ 11 Fibration exact sequence No
Remark 6.1 (Scope of Applications).
While this paper focuses on SVK, our formalization includes other fundamental group techniques:
  • Product formula: $\pi_{1}(A\times B)\simeq\pi_{1}(A)\times\pi_{1}(B)$ (used for torus)
  • Fibration sequences: $\pi_{2}(B)\to\pi_{1}(F)\to\pi_{1}(E)\to\pi_{1}(B)$ (used for $\mathbb{CP}^{n}$)
  • Direct encode-decode: For HITs with explicit path constructors (Circle, Klein bottle)
The torus and $\mathbb{CP}^{n}$ are included to show the breadth of the formalization, though they do not use SVK.
6.1. The Wedge Sum: Free Product of Fundamental Groups
Theorem 6.2 (Fundamental Group of Wedge Sum).
For pointed types $(A,a_{0})$ and $(B,b_{0})$:
$\pi_{1}(A\vee B)\simeq\pi_{1}(A)\ast\pi_{1}(B)$
Proof.
The wedge sum is the pushout $\mathrm{Pushout}(A,B,\mathbf{1})$. Since $\pi_{1}(\mathbf{1})=1$, the amalgamation is trivial, yielding the free product. ∎
In Lean:
1noncomputable def wedgeFundamentalGroupEquiv
2 [HasWedgeSVKDecodeBijective a0 b0] :
3 SimpleEquiv (pi_1(Wedge A B a0 b0, wedgeBase))
4 (FreeProduct (pi_1(A, a0)) (pi_1(B, b0)))
6.2. The Figure-Eight Space
Theorem 6.3 (Fundamental Group of Figure-Eight).
$\pi_{1}(S^{1}\vee S^{1})\simeq\mathbb{Z}\ast\mathbb{Z}$
The figure-eight has a non-abelian fundamental group:
1def wordAB : FreeProductWord Int Int := .consLeft 1 (.consRight 1 .nil)
2def wordBA : FreeProductWord Int Int := .consRight 1 (.consLeft 1 .nil)
3
4theorem wordAB_ne_wordBA : wordAB != wordBA := by
5 intro h; cases h -- constructors are distinct
Remark 6.4 (Non-Commutativity).
The fact that $ab\neq ba$ in $\mathbb{Z}\ast\mathbb{Z}$ demonstrates that $\pi_{1}(S^{1}\vee S^{1})$ is non-abelian. This is a key distinguishing feature from the torus, where $\pi_{1}(T^{2})\cong\mathbb{Z}\times\mathbb{Z}$ is abelian.
6.3. Spheres
Theorem 6.5 (Fundamental Group of $n$-Sphere).
For $n\geq 2$:
$\pi_{1}(S^{n})\simeq 1$
Proof.
The $n$-sphere is the (n1)(n-1)-fold iterated suspension: $S^{n}=\Sigma^{n-1}(S^{1})$. By SVK applied to suspension:
$\pi_{1}(\Sigma A)\simeq 1\ast_{\pi_{1}(A)}1=1$
since both “north” and “south” components are points with trivial $\pi_{1}$. ∎
6.4. The Torus (Product Formula, Not SVK)
Theorem 6.6 (Fundamental Group of Torus).
$\pi_{1}(T^{2})\simeq\mathbb{Z}\times\mathbb{Z}$
Remark 6.7 (Torus Construction—Not SVK).
The torus does not use SVK. Instead, we construct $T^{2}:=S^{1}\times S^{1}$ and apply the product formula:
$\pi_{1}(A\times B,(a_{0},b_{0}))\simeq\pi_{1}(A,a_{0})\times\pi_{1}(B,b_{0})$
This requires no additional axioms beyond Circle. The product formula is proved directly from the definition of paths in product types:
$\mathrm{Path}((a_{1},b_{1}),(a_{2},b_{2}))\simeq\mathrm{Path}(a_{1},a_{2})\times\mathrm{Path}(b_{1},b_{2})$
We include the torus here to demonstrate that the formalization supports multiple $\pi_{1}$ calculation techniques, not just SVK.
6.5. The Klein Bottle
Theorem 6.8 (Fundamental Group of Klein Bottle).
$\pi_{1}(K)\simeq\mathbb{Z}\rtimes\mathbb{Z}=\langle a,b\mid aba^{-1}b=1\rangle$
where the semidirect product has $\mathbb{Z}$ acting on $\mathbb{Z}$ by negation.
In Lean:
1structure KleinBottlePresentation where
2 -- Elements are pairs (m, n) where:
3 -- m represents powers of the loop that reverses orientation
4 -- n represents powers of the loop that preserves orientation
5 m : Int
6 n : Int
7
[…]
9-- (m1, n1) * (m2, n2) = (m1 + m2, (-1)^m2 * n1 + n2)
6.6. The Real Projective Plane
Theorem 6.9 (Fundamental Group of Projective Plane).
$\pi_{1}(\mathbb{RP}^{2})\simeq\mathbb{Z}_{2}=\langle a\mid a^{2}=1\rangle$
Proof.
$\mathbb{RP}^{2}$ can be constructed by attaching a disk to a circle via the map $z\mapsto z^{2}$. The fundamental group has one generator (the loop) with the relation that going around twice is contractible. ∎
6.7. Orientable Surfaces of Genus $g$
Theorem 6.10 (Fundamental Group of Orientable Surface).
$\pi_{1}(\Sigma_{g})\simeq\langle a_{1},b_{1},\ldots,a_{g},b_{g}\mid[a_{1},b_{1}]\cdots[a_{g},b_{g}]=1\rangle$
where $a,b]=aba^{-1}b^{-1}$ is the commutator.
In Lean:
1def OrientableSurfaceWord (g : Nat) : Type :=
2 List (Fin (2 * g) * Bool) -- generator index and sign
3
4inductive SurfaceRelation (g : Nat) :
5 OrientableSurfaceWord g -> OrientableSurfaceWord g -> Prop
6 | commutator_product :
7 -- [a_1, b_1] ... [a_g, b_g] = 1
8 SurfaceRelation g (commutatorProduct g) []
6.8. Non-Orientable Surfaces
Theorem 6.11 (Fundamental Group of Non-Orientable Surface).
$\pi_{1}(N_{n})\simeq\langle a_{1},\ldots,a_{n}\mid a_{1}^{2}\cdots a_{n}^{2}=1\rangle$
Example 6.12 (Special Cases).
  • $N_{1}=\mathbb{RP}^{2}$: $\pi_{1}(N_{1})=\langle a\mid a^{2}=1\rangle\cong\mathbb{Z}_{2}$
  • $N_{2}=K$ (Klein bottle): $\pi_{1}(N_{2})=\langle a,b\mid a^{2}b^{2}=1\rangle\cong\mathbb{Z}\rtimes\mathbb{Z}$
6.9. Lens Spaces
Theorem 6.13 (Fundamental Group of Lens Space).
For coprime $p,q$:
$\pi_{1}(L(p,q))\simeq\mathbb{Z}/p\mathbb{Z}$
Remark 6.14 (Lens Space Construction).
Lens spaces $L(p,q)$ are quotients of $S^{3}$ by cyclic group actions. The fundamental group depends only on $p$, not $q$; however, the spaces $L(p,q)$ and $L(p,q^{\prime})$ may be non-homeomorphic for different values of $q$.
6.10. Bouquet of $n$ Circles
Theorem 6.15 (Fundamental Group of $n$-Bouquet).
$\pi_{1}(\bigvee_{n}S^{1})\simeq F_{n}$
where $F_{n}$ is the free group on $n$ generators.
7. Extended Features
7.1. Higher Homotopy Groups
Definition 7.1 (Higher Homotopy Groups).
The $n$-th homotopy group is defined via iterated loop spaces:
$\pi_{n}(A,a):=\pi_{1}(\Omega^{n-1}(A,a),\rho^{n-1})$
where $\Omega^{n}$ denotes $n$-fold iteration of the loop space.
Theorem 7.2 (Eckmann-Hilton).
For $n\geq 2$, $\pi_{n}(A,a)$ is abelian.
Proof.
The Eckmann-Hilton argument shows that when two binary operations share a common unit and satisfy an interchange law, they must coincide and be commutative. In $\Omega^{2}(A,a)$, both horizontal and vertical composition of 2-loops satisfy these conditions. ∎
In Lean:
1def Loop2Space (A : Type u) (a : A) : Type u :=
2 Derivation_2 (Path.refl a) (Path.refl a)
3
4def PiTwo (A : Type u) (a : A) : Type u :=
5 Quotient (Loop2Setoid A a)
6
7theorem piTwo_comm (x y : PiTwo A a) : PiTwo.mul x y = PiTwo.mul y x := by
8 induction x using Quotient.ind
9 induction y using Quotient.ind
10 apply Quotient.sound
11 exact contractibility_3 _ _
7.2. The Weak $\omega$-Groupoid Structure
Theorem 7.3 (Types are Weak $\omega$-Groupoids).
Via computational paths, every type carries the structure of a weak $\omega$-groupoid:
  • 0-cells: points of AA
  • 1-cells: paths $p:\mathrm{Path}(a,b)$
  • 2-cells: derivations $\mathcal{D}_{2}(p,q)$ witnessing $p\sim q$
  • 3-cells: meta-derivations $\mathcal{D}_{3}(\alpha,\beta)$
  • $n$-cells: $\mathcal{D}_{n}$ for all $n$
In Lean:
1-- 2-cells: derivations between paths
2inductive Derivation_2 : Path a b -> Path a b -> Type u
3 | refl (p : Path a b) : Derivation_2 p p
4 | step (s : Step p q) : Derivation_2 p q
5 | inv (d : Derivation_2 p q) : Derivation_2 q p
6 | vcomp (d1 : Derivation_2 p q) (d2 : Derivation_2 q r) : Derivation_2 p r
7
8-- 3-cells: derivations between 2-cells
9inductive Derivation_3 : Derivation_2 p q -> Derivation_2 p q -> Type u
10 | refl (d : Derivation_2 p q) : Derivation_3 d d
11 | step (m : MetaStep_2 d1 d2) : Derivation_3 d1 d2
12 | inv (m : Derivation_3 d1 d2) : Derivation_3 d2 d1
13 | vcomp (m1 : Derivation_3 d1 d2) (m2 : Derivation_3 d2 d3) : Derivation_3 d1 d3
7.3. Truncation Levels
Definition 7.4 (Truncation Levels).
Following HoTT, we define truncation levels:
  • IsContr(AA): AA is contractible (has a unique point)
  • IsProp(AA): any two points are equal (“propositions”)
  • IsSet(AA): any two parallel paths are equal (“sets”)
  • IsGroupoid(AA): any two parallel 2-cells are equal
In Lean:
1class IsContr (A : Type u) where
2 center : A
3 contr : forall a, Path center a
4
5class IsProp (A : Type u) where
6 allEq : forall (a b : A), Path a b
7
8class IsSet (A : Type u) where
9 pathEq : forall (a b : A) (p q : Path a b), RwEq p q
10
11class IsGroupoid (A : Type u) where
12 derivEq : forall (a b : A) (p q : Path a b) (d e : Derivation_2 p q),
13 Derivation_3 d e
7.4. Covering Space Theory
Definition 7.5 (Covering Space).
A covering space of $B$ with fiber $F$ consists of:
  • A total space $E$
  • A projection $p:E\to B$
  • A $\pi_{1}(B)$-action on $F$
  • Fiber equivalence: $p^{-1}(b)\simeq F$ for each $b$
In Lean:
1structure CoveringSpace (B : Type u) (F : Type u) (b0 : B) where
2 totalSpace : Type u
3 proj : totalSpace -> B
4 fiber_equiv : forall b, Equiv (Fiber proj b) F
5 deck_action : pi_1(B, b0) -> Equiv F F
6 action_coherent : -- action is compatible with path lifting
7.5. Fibration Theory
Definition 7.6 (Fibration).
A fibration $p:E\to B$ with fiber $F$ admits a long exact sequence:
$\cdots\to\pi_{2}(F)\to\pi_{2}(E)\to\pi_{2}(B)\to\pi_{1}(F)\to\pi_{1}(E)\to\pi_{1}(B)\to 1$
7.6. Eilenberg-MacLane Spaces
Theorem 7.7 ($S^{1}$ as $K(\mathbb{Z},1)$).
The circle is the Eilenberg-MacLane space $K(\mathbb{Z},1)$:
  • $\pi_{1}(S^{1})\cong\mathbb{Z}$
  • $\pi_{n}(S^{1})\cong 1$ for $n\geq 2$
7.7. Path Simplification Tactics
The formalization includes automated tactics for path reasoning:
1-- Basic simplification using groupoid laws
2path_simp -- refl . p ~ p, p . refl ~ p, etc.
3
4-- Full automation (~25 simp lemmas)
5path_auto -- handles most standalone RwEq goals
6
7-- Specialized tactics
8path_rfl -- close p ~ p
9path_canon -- normalize to canonical form
10path_cancel_left -- p^-1 . p ~ refl
11path_cancel_right -- p . p^-1 ~ refl
12path_congr_left h -- apply hypothesis on right
13path_congr_right h -- apply hypothesis on left
8. The Lean 4 Formalization
8.1. Why Axioms Are Necessary
Lean 4 does not natively support higher-inductive types (HITs). Unlike Agda’s --cubical mode or Coq’s HoTT library, there is no built-in mechanism for defining types with both point and path constructors. We therefore axiomatize HITs, introducing axioms only where strictly necessary.
Definition 8.1 (Kernel Axiom).
A kernel axiom is a Lean axiom declaration that extends the trusted kernel. These are used sparingly for:
  1. (1)
    HIT type declarations: The type itself (e.g., Circle : Type u)
  2. (2)
    Point constructors: Base points (e.g., circleBase : Circle)
  3. (3)
    Path constructors: Generating paths (e.g., circleLoop : Path circleBase circleBase)
  4. (4)
    Higher path constructors: 2-cells, surface relations
  5. (5)
    Recursion principles: Eliminators with $\beta$-rules
Definition 8.2 (Typeclass Assumption).
A typeclass assumption is a hypothesis provided via Lean’s typeclass mechanism. These are not kernel axioms—they are parameters that users instantiate. We use these for:
  1. (1)
    Computation rules: HasRecGlueRwEq (how rec acts on glue)
  2. (2)
    Naturality: HasGlueNaturalLoopRwEq (glue commutes with paths)
  3. (3)
    Encode/decode data: HasPushoutSVKEncodeQuot, etc.
Remark 8.3 (Axioms vs Assumptions).
The distinction is crucial:
  • Kernel axioms (36 total) are permanently trusted by Lean. We minimize these.
  • Typeclass assumptions are user-provided hypotheses. They appear in theorem statements as [HasFoo] parameters, making dependencies explicit.
For example, seifertVanKampenEquiv requires several typeclass assumptions but introduces no new kernel axioms—users must provide evidence for the assumptions.
Remark 8.4 (Consistency of the Encode/Decode Axioms).
A natural question arises: why doesn’t the fundamental group collapse? In Lean’s proof-irrelevant Prop, all equality proofs between the same terms are equal. If $\pi_{1}$ depended only on equality proofs, all loops would be identified.
The key to consistency is threefold:
  1. (1)
    Explicit step lists (see Remark 2.1): Paths store explicit steps : List (Step A), and $\sim$ identifies paths only when their step sequences can be transformed via the Step rules—not merely when their underlying equality proofs coincide.
  2. (2)
    No path collapse rule: The Step relation contains no rule that identifies arbitrary paths with the same endpoints. In particular, there is no “canonicalization” rule $p\rightsquigarrow\mathrm{ofEq}(p.\mathrm{proof})$ that would reduce any path to a canonical form based solely on its equality proof.
  3. (3)
    Axiomatized classification data: The encode/decode equivalences (e.g., $\pi_{1}(S^{1})\simeq\mathbb{Z}$) are assumed via typeclass assumptions, not derived. We do not assert that transport along a self-equality computes nontrivially—such an assertion would be inconsistent with Lean’s proof-irrelevant Prop. Instead, we assume the classification result: there exists an equivalence between $\pi_{1}$ and the expected group, with the round-trip properties holding.
This design follows the HoTT tradition of axiomatizing what cannot be computed in the ambient type theory. The typeclass assumptions for encode/decode are analogous to Licata-Shulman’s encode-decode method [10], but formulated as external assumptions rather than internal proofs.
8.2. Justifying the 36 Kernel Axioms
Each HIT requires axioms because Lean cannot construct them from primitive types:
HIT Why Axioms Are Needed
Circle Needs a non-trivial loop $\mathrm{loop}:\mathrm{Path}(\mathrm{base},\mathrm{base})$ that is not $\rho$. No Lean type has this property without axioms.
Sphere Needs 2-dimensional structure (meridians collapsing at poles).
Klein bottle Needs the non-orientable surface relation $aba^{-1}b=1$.
$\mathbb{RP}^{2}$ Needs $a^{2}=1$ as a path relation.
$\Sigma_{g}$, $N_{n}$ Need genus-dependent relations on generators.
Cylinder Needs both a loop and a segment with a 2-cell between them.
Remark 8.5 (Axiom-Free Constructions).
Some spaces require no additional kernel axioms beyond those already listed:
  • Torus: Defined as $T^{2}:=S^{1}\times S^{1}$ (uses Circle axioms only)
  • Möbius band: Homotopy equivalent to $S^{1}$ (uses Circle axioms only)
  • Wedge sum: Defined as $\mathrm{Pushout}(A,B,\mathbf{1})$ (zero kernel axioms—Pushout is Quot-based)
  • Suspension: Defined as $\mathrm{Pushout}(\mathbf{1},\mathbf{1},A)$ (zero kernel axioms—Pushout is Quot-based)
  • Sphere: Defined as $S^{n}:=\Sigma^{n}(\mathbf{1})$ via iterated suspension (zero kernel axioms)
8.3. When to Use Each SVK Assumption
The split typeclass architecture allows users to provide minimal hypotheses:
Assumption When to Use
HasGlueNaturalLoopRwEq Always required for SVK. Automatically derived when $C$ is a subsingleton (wedge) or when $A$, $B$ satisfy Axiom K.
HasPushoutSVKEncodeQuot When you have an explicit encode function.
HasPushoutSVKDecodeEncode Standard encode-decode round-trip.
HasPushoutSVKEncodeDecode Round-trip up to amalgamation equivalence.
HasPushoutSVKDecodeAmalgBijective Choice-based interface when you only know decode is bijective (Prop-level), without explicit encode.
8.4. Architecture
The Lean 4 formalization is organized as follows:
Module Content
Path/Basic/Core.lean Path type, refl, symm, trans
Path/Basic/Congruence.lean congrArg, transport
Path/Rewrite/Step.lean Single-step rewrites (50+ rules)
Path/Rewrite/RwEq.lean Rewrite equality
Path/Rewrite/PathTactic.lean Automated tactics
Path/Homotopy/Loops.lean Loop spaces
Path/Homotopy/FundamentalGroup.lean $\pi_{1}$ definition
Path/Homotopy/HigherHomotopy.lean $\pi_{n}$ for $n\geq 2$
Path/HIT/Circle.lean Circle HIT
Path/HIT/Pushout.lean Pushout HIT
Path/HIT/PushoutPaths.lean SVK theorem
Path/HIT/FigureEight.lean $S^{1}\vee S^{1}$
Path/HIT/OrientableSurface.lean $\Sigma_{g}$
Path/HIT/NonOrientableSurface.lean $N_{n}$
Path/HIT/KleinBottle.lean Klein bottle
Path/HIT/ProjectivePlane.lean $\mathbb{RP}^{2}$
Path/HIT/LensSpace.lean $L(p,q)$
Path/OmegaGroupoid.lean Weak $\omega$-groupoid structure
8.5. Kernel Axiom Inventory
The formalization uses exactly 36 kernel axioms, all for higher-inductive types:
Table 2. Kernel axioms by HIT
HIT Axioms Description
Circle 7 Type, base, loop, rec, rec_base, rec_loop, ind
[…]
KleinBottle 5 Type, base, loopA, loopB, surface
OrientableSurface 5 Type, base, loopA, loopB, surfaceRelation
NonOrientableSurface 4 Type, base, loop, 2-cell
ProjectivePlane 4 Type, base, loop, loopSquare
Total 36
Notable non-axioms (constructions that add 0 kernel axioms):
  • Pushout: Defined as Quot (PushoutRel A B C f g) using Lean’s built-in quotient type—no kernel axioms. Constructors inl, inr, glue are definitions, not axioms.
  • Sphere: Defined as $S^{n}:=\Sigma^{n}(\mathbf{1})$ via suspension of the unit type—uses Pushout.
  • Torus: Constructed as $S^{1}\times S^{1}$—uses only Circle axioms.
  • Möbius band: Homotopy equivalent to $S^{1}$—uses only Circle axioms.
  • Wedge, Suspension: Defined as special cases of Pushout.
  • SVK encode/decode: Typeclass assumptions, not kernel axioms.
  • Lens spaces, Bouquet: Constructed from existing HITs.
Remark 8.6 (Pushout Design).
Unlike HITs such as Circle or Klein bottle (which require kernel axioms for their path constructors), the Pushout is implemented using Lean’s built-in Quot type. The constructors inl, inr are Quot.mk, and glue is Path.ofEq (Quot.sound ...).
The computation rules (how rec acts on glue, glue naturality) are provided via typeclass assumptions (HasRecGlueRwEq, HasGlueNaturalLoopRwEq), not kernel axioms. This design means Pushout adds zero kernel axioms; all structure comes from user-instantiated typeclasses.
8.6. Opt-in Axiom Imports
For convenience, opt-in imports provide assumption-free APIs:
1-- Import to get circlePiOneEquivInt’ with no hypotheses
2import ComputationalPaths.Path.HIT.CirclePiOneAxiom
3
4-- Import to get torusPiOneEquivIntProd’ with no hypotheses
5import ComputationalPaths.Path.HIT.TorusPiOneAxiom
6
7-- Import to get wedgeFundamentalGroupEquiv’ with no hypotheses
8import ComputationalPaths.Path.HIT.WedgeSVKAxiom
8.7. Statistics
Metric Value
Total lines of Lean 41,130
Number of modules 107
Kernel axioms 36
Theorems/lemmas 1,200+
Definitions 500+
8.8. Assumption Manifest
For reviewer clarity, we provide a unified summary of all assumptions in the formalization.
Part A: Kernel Axioms (36 total, all for HIT type/constructor declarations)
HIT Kernel Axioms (Lean names)
Circle (7) Circle, circleBase, circleLoop, circleRec, circleRec_base, circleRec_loop, circleInd
Cylinder (11) Cylinder, cylinderBase0/1, cylinderSeg, cylinderLoop0/1, cylinderSurf, cylinderRec, cylinderRec_base0/1, cylinderRec_loop0
KleinBottle (5) KleinBottle, kleinBase, kleinLoopA/B, kleinSurf
OrientableSurface (5) OrientableSurface, base, loopA, loopB, surfaceRelation
NonOrientableSurface (4) NonOrientableSurface, nonOrientableBase, nonOrientableLoop, nonOrientableSurface2Cell
ProjectivePlane (4) ProjectivePlane, projectiveBase, projectiveLoop, projectiveLoopSquare
Not kernel axioms: Pushout (Quot-based), Sphere (suspension), Torus ($S^{1}\times S^{1}$), Wedge, Suspension, Lens spaces, Bouquet.
Part B: Typeclass Assumptions for SVK (Theorem 5.1)
Typeclass Role Status
HasGlueNaturalLoopRwEq Glue naturality Required; auto-derived for subsingletons
HasPushoutSVKEncodeQuot Encode map exists Assumed (user-provided)
HasPushoutSVKDecodeEncode Decode-encode round-trip Assumed (user-provided)
HasPushoutSVKEncodeDecode Encode-decode round-trip Assumed (user-provided)
HasPushoutSVKDecodeAmalgBijective Alternative: bijectivity only Assumed (choice-based)
Proved unconditionally: Decode map construction, decode respects amalgamation (given glue naturality).
Part C: Opt-in Axiom Instances (for assumption-free APIs)
These are additional kernel axioms that globally instantiate typeclasses:
  • instHasCirclePiOneEncodeAxiom in CirclePiOneAxiom.lean
  • instHasKleinPiOneEncodeAxiom in KleinPiOneAxiom.lean
  • instHasProjectivePiOneEncodeAxiom in ProjectivePiOneAxiom.lean
  • instHasLensPiOneEncodeAxiom in LensPiOneAxiom.lean
  • instHasWedgeSVKDecodeBijectiveAxiom in WedgeSVKAxiom.lean
These are not counted in the 36 core axioms; they are optional imports for convenience.
9. Library Highlights
Beyond the SVK framework, the library includes extended features for advanced homotopy-theoretic constructions:
9.1. Higher Homotopy Groups
The higher homotopy groups $\pi_{n}(A,a)$ for $n\geq 2$ are defined via iterated loop spaces:
$\pi_{n}(A,a):=\pi_{1}(\Omega^{n-1}(A,a),\rho)$
Key results include:
  • Abelianness of $\pi_{2}$ via Eckmann-Hilton (HigherHomotopy.lean)
  • Long exact sequence of a fibration (Fibration.lean)
9.2. Weak $\omega$-Groupoid Structure
Types are shown to form weak $\omega$-groupoids via the derivation tower:
  • 1-cells: paths $\mathrm{Path}(a,b)$
  • 2-cells: derivations $\mathcal{D}_{2}(p,q)$ witnessing $p\sim q$
  • 3-cells: derivations $\mathcal{D}_{3}(d_{1},d_{2})$ between derivations
  • And so on, with truncation at level 3 (OmegaGroupoid.lean)
9.3. Truncation Levels
The library formalizes HoTT-style truncation levels:
  • IsContr: Contractible types (all points equal)
  • IsProp: Propositions (all proofs equal)
  • IsSet: Sets (all paths equal)
  • IsGroupoid: Groupoids (all 2-paths equal)
Module: Truncation.lean
9.4. Covering Spaces
Covering space theory with $\pi_{1}$-actions:
  • Fiber transport action (CoveringSpace.lean)
  • Galois correspondence framework (CoveringClassification.lean)
9.5. Additional Features
  • Eilenberg-MacLane spaces: $S^{1}=K(\mathbb{Z},1)$ characterization (EilenbergMacLane.lean)
  • Mayer-Vietoris: Abelianized SVK sequence (MayerVietoris.lean)
  • Path tactics: Automation via path_simp, path_auto, path_normalize (PathTactic.lean)
10. Conclusion
We have presented a modular SVK framework within the computational paths setting, demonstrating that:
  1. (1)
    Pushouts can be implemented via Lean’s quotient types with zero kernel axioms, while computation rules are supplied as typeclass assumptions.
  2. (2)
    The encode-decode method works smoothly with split assumptions, supporting both data-level and Prop-level (choice-based) interfaces. The SVK equivalence is parametric in user-supplied encode/decode structure.
  3. (3)
    The framework yields instantiations for all closed surfaces, lens spaces, projective spaces, and more—with explicit dependency tracking via typeclasses.
  4. (4)
    The entire development comprises 41,130 lines of Lean 4 code with only 36 kernel axioms (for HIT type/constructor declarations).
10.1. Comparison with HoTT Approaches
Table 3. Comparison with Standard HoTT
Aspect Standard HoTT Computational Paths
Path equality Identity type ($\mathrm{Id}$) Syntactic ($\sim$)
Coherence Abstract existence Explicit rewrite derivations
SVK assumptions Monolithic Split/modular typeclasses
Encode function Code family via $J$ Typeclass assumption
$\omega$-groupoid Higher identity types Derivation tower
Decidable under termination + confluence; termination not mechanized here (see §2.4).
10.2. Limitations
  1. (1)
    Axiom usage: The encode/decode equivalences are axiomatized via typeclasses rather than proved. In cubical Agda, where transport computes, these would be provable. This is an inherent limitation of standard Lean’s proof-irrelevant Prop—we cannot prove that transport along a non-trivial loop acts non-trivially. See Remark 8.4 for why this design is nonetheless consistent.
  2. (2)
    Explicit step storage: Paths store explicit List (Step A) to distinguish loops—a departure from the “thin” identity types of standard MLTT. This is essential for non-trivial $\pi_{1}$ (see Remark 2.1).
  3. (3)
    Path-connectedness assumption: The SVK theorem requires path-connected types.
10.3. Future Work
  1. (1)
    Cubical type theory comparison: Relate computational paths to cubical approaches.
  2. (2)
    Fully constructive encode: Implement code families to eliminate encode axioms.
  3. (3)
    Covering space classification: Prove the Galois correspondence between covering spaces and subgroups of $\pi_{1}$.
  4. (4)
    Higher Hopf fibrations: Quaternionic ($S^{3}\to S^{7}\to S^{4}$) and octonionic.
  5. (5)
    3-manifold classification: Seifert fibered spaces, graph manifolds.
  6. (6)
    Homology theory: Extend from $\pi_{1}$ to full homology groups.
Appendix A Application Inventory
The following table provides the exact Lean theorem names and module paths for the headline results. All theorems depend on typeclass assumptions as indicated; “opt-in” means an axiom import is available for assumption-free use.
Result Lean Name Module Deps
$\pi_{1}(S^{1})\simeq\mathbb{Z}$ circlePiOneEquivInt CircleStep Opt-in
$\pi_{1}(T^{2})\simeq\mathbb{Z}^{2}$ torusPiOneEquivIntProd TorusStep Opt-in
$\pi_{1}(S^{2})\simeq 1$ sphere2PiOneEquivUnit Sphere None
$\pi_{1}(S^{1}\vee S^{1})$ figureEightPiOneEquiv FigureEight SVK
SVK schema seifertVanKampenEquiv PushoutPaths SVK
SVK choice svkEquiv_of_bijective PushoutPaths Bij.
$\pi_{1}(K)$ kleinPiOneEquivSemidirect KleinBottleStep Opt-in
$\pi_{1}(\mathbb{RP}^{2})\simeq\mathbb{Z}_{2}$ projectivePiOneEquivZ2 ProjectivePlane Opt-in
$\pi_{1}(\Sigma_{g})$ piOneEquivPresentation OrientableSurface SVK
$\pi_{1}(N_{n})$ piOneEquivPresentation NonOrientableSurface SVK
$\pi_{1}(L(p,q))$ lensPiOneEquivZMod LensSpace Opt-in
Wedge wedgeFundamentalGroupEquiv WedgePaths Opt-in
Legend: Opt-in: axiom import provides global instance. SVK: requires encode/decode typeclasses. Bij.: requires HasPushoutSVKDecodeAmalgBijective. None: no extra assumptions.